### v1

# Rings of Hilbert modular forms, computations on Hilbert modular surfaces, and the Oda-Hamahata conjecture

Abstract.
The modularity of an elliptic curve $E/{\mathbb{Q}}$ can be expressed either as an analytic statement that the $L$-function is the Mellin transform of a modular form, or as a geometric statement that $E$ is a quotient of a modular curve $X_{0}(N)$. For elliptic curves over number fields these notions diverge; a conjecture of Hamahata asserts that for every elliptic curve $E$ over a totally real number field there is a correspondence between a Hilbert modular variety and the product of the conjugates of $E$. In this paper we prove the conjecture by explicit computation for many cases where $E$ is defined over a real quadratic field and the geometric genus of the Hilbert modular variety is $1$.
Key words and phrases:
Hilbert modular surfaces, elliptic curves, real quadratic fields, modularity
2020 Mathematics Subject Classification:
14G35, 14Q10, 11G05, 11F41, 14Q25
1. Introduction
One of the most fundamental discoveries in the arithmetic of elliptic curves over ${\mathbb{Q}}$ is the modularity theorem, previously the Shimura-Taniyama-Weil conjecture and later proved in [6], building on work of Wiles, Taylor-Wiles, and many other mathematicians. We may state it briefly in two ways:
Theorem 1.1.
Let $E$ be an elliptic curve over ${\mathbb{Q}}$ and let $L(s,E)$ be its $L$-function as defined in [33, Ex. 8.19]. Let $N=N_{E}$ be the conductor of $E$ and let $X_{0}(N)$ be the modular curve parametrizing elliptic curves with a cyclic subgroup of order $N$. Then:
  1. (analytic)
    $L(s,E)$ is the Mellin transform of a Hecke eigenform of weight $2$ and level $N$. (Concretely, the coefficients of the Dirichlet series of $L$ are the same as those of the Fourier expansion of $f$.)
  2. (geometric)
    There is a dominant map $X_{0}(N)\to E$.
Our main goal in this paper is to prove a few special cases of a form of geometric modularity for elliptic curves over real quadratic fields. We will state the result and then devote the rest of the introduction to explaining the statement.
Theorem 1.2.
Let $E$ be an elliptic curve over a real quadratic field belonging to an isogeny class with one of the following LMFDB labels or their Galois conjugates: 2.2.5.1/31.1-a, 2.2.5.1/55.1-a, 2.2.5.1/80.1-a, 2.2.8.1/9.1-a, 2.2.8.1/17.1-a, 2.2.8.1/32.1-a, 2.2.13.1/9.1-a, 2.2.17.1/4.1-a, 2.2.17.1/8.3-a. Then there is a correspondence (not necessarily defined over ${\mathbb{Q}}$) between the Hilbert modular variety of level equal to the conductor of $E$ and the Kummer surface of $E\times E^{\sigma}$.
To give some context for this result, let us more generally consider an elliptic curve $E$ over a totally real number field $K$. In order to describe the situation we begin with a brief explanation of the concept of a Hilbert modular form.
Definition 1.3.
[36, I.1, II.7] Let $K$ be a totally real number field of degree $d$ over ${\mathbb{Q}}$, let $\mathcal{H}$ be the upper half-plane, and let $\mathcal{H}_{\pm}$ be the union of the upper and lower half-planes. Then $\operatorname{GL}_{2}^{+}({\mathbb{R}})$ acts on $\mathcal{H}$ and so $\operatorname{GL}_{2}^{+}({\mathbb{R}})^{d}$ acts on $\mathcal{H}^{d}$. Choosing an ordering $\rho_{1},\dots,\rho_{d}$ of the real embeddings of $K$, we obtain an embedding $\operatorname{GL}_{2}^{+}(K)\hookrightarrow\operatorname{GL}_{2}^{+}({\mathbb{R}})^{d}$ and hence an action of $\operatorname{GL}_{2}^{+}(K)$ on $\mathcal{H}^{d}$.
Definition 1.4.
Let $I\subseteq{\mathcal{O}}_{K}$ be a nonzero ideal. Then $\Gamma_{0}(I)$ is the subgroup of $\operatorname{SL}_{2}({\mathcal{O}}_{K})$ of matrices whose lower left entry belongs to $I$. (One also adapts the usual definitions for $\operatorname{SL}_{2}({\mathbb{Z}})$ to define $\Gamma_{1}(I),\Gamma(I)$, but these will not be used in this paper.)
We remark [36, p. 6, p. 15] that $\operatorname{SL}_{2}({\mathcal{O}}_{K})$ acts properly discontinuously and with finite stabilizers on $\mathcal{H}^{d}$. This is the reason to act on $\mathcal{H}^{d}$; if we only used some of the embeddings to define the action, then every orbit would be dense and the quotient would not have a reasonable topology.
Definition 1.5.
[36, Definition I.6.1] Let $d>1$ and let $w_{1},\dots,w_{d}$ be integers. A Hilbert modular form of weights $(w_{1},\dots,w_{d})$ and level $I$ is a holomorphic function $f$ on $\mathcal{H}^{d}$ such that, for every $M=\begin{pmatrix}p&q\\ r&s\end{pmatrix}\in\Gamma_{0}(I)$ and $(z_{1},\dots,z_{d})\in\mathcal{H}^{d}$, we have
$f(z_{1},\dots,z_{d})=\prod_{i=1}^{d}(\rho_{i}(r)z_{i}+\rho_{i}(s))^{w_{i}}f\left(\frac{\rho_{1}(p)z_{1}+\rho_{1}(q)}{\rho_{1}(r)z_{1}+\rho_{1}(s)}\right),\dots,f\left(\frac{\rho_{d}(p)z_{1}+\rho_{d}(q)}{\rho_{d}(r)z_{1}+\rho_{d}(s)}\right).$
We remark that Hilbert modular forms constitute a ${\mathbb{Z}}^{d}$-graded ring, the product of forms of weights $(w_{1},\dots,w_{d})$ and $(w_{1}^{\prime},\dots,w_{d}^{\prime})$ being a form of weight $(w_{1}+w_{1}^{\prime},\dots,w_{d}+w_{d}^{\prime})$. As with classical modular forms, Hilbert modular forms have a sort of Fourier expansion with coefficients indexed by totally positive elements of the inverse different of ${\mathcal{O}}_{K}$. For certain purposes it is better to consider the action of a slightly larger group, such as a subgroup of $\operatorname{GL}_{2}({\mathcal{O}}_{K})$, on the product of $d$ copies of $\mathcal{H}_{\pm}$, the union of the upper and lower half-planes. For this paper it is mostly not necessary to do this and so we leave the minor modifications necessary to the reader.
We now return to elliptic curves over $K$. Let $E$ be such a curve and let its conductor be $I$. To generalize the algebraic statement of modularity is fairly straightforward: one requires that the coefficients of the $L$-function of $E$ match the Hecke eigenvalues of a Hilbert modular form $f$. This is known for all elliptic curves over real quadratic fields and many over fields of higher degree: see [10, Theorems 1-2].
To state a reasonable geometric modularity conjecture requires some thought. Before doing so, we introduce the Hilbert modular variety associated to a group $\Gamma$. For simplicity we will only do this over ${\mathbb{C}}$ by a complex-analytic construction.
Definition 1.6.
Let $\Gamma$ be a subgroup of $\operatorname{GL}_{2}^{+}(K)$ commensurable with $\operatorname{SL}_{2}({\mathcal{O}}_{K})$. Then the Hilbert modular variety $\Gamma\backslash\mathcal{H}^{d}$ is the quotient of $\mathcal{H}^{d}$ by $\Gamma$ in the category of complex analytic spaces. If $\Gamma=\Gamma_{0}(I)$, then we write $Y_{0}(I)$ for the quotient.
From this point of view it is not at all obvious that $\Gamma\backslash\mathcal{H}^{d}$ has an algebraic structure, much less that it can be defined over a number field, but this is true [36, Thm. IV.7.1].
Proposition 1.7.
Let $J_{1},\dots,J_{h_{I}}$ be ideals representing the cosets of the group of fractional ideals of ${\mathcal{O}}_{K}$ modulo those generated by an element that is $1\bmod I$. Then there exists a projective variety that is constructed from $Y_{0}(I)$ by adding points $P_{1},\dots,P_{h_{I}}$ in canonical bijection with the $J_{i}$.
This variety, which we will denote $H_{I,K}$, is the famous Baily-Borel compactification [36, II.7], which is to say that it is Proj of the ring of Hilbert modular forms defined above.
Remark 1.8.
In contrast to the classical case where the presence of $-I_{2}$ in $\Gamma_{0}(N)$ implies that there are no nonzero cusp forms of odd weight, such forms may exist here. However, we will always consider Proj of the subring of Hilbert modular forms of even weight. This is the second Veronese embedding and is therefore isomorphic to Proj of the full ring by [34, Exercise 7.4.D].
What is the geometric relation between $H_{I,K}$ and an elliptic curve $E$ over $K$ of conductor $I$? It is quite unreasonable to expect a nonconstant map $H_{I,K}\to E$ as in the classical case, since $H_{I,K}$ is defined over ${\mathbb{Q}}$ and $E$ only over a number field. We might replace $E$ by the product of its conjugates, which can be defined over ${\mathbb{Q}}$, but this will still not admit any nonconstant maps from $H_{I,K}$. The reason is that $h^{1}(H_{I,K})=0$ for $d>1$ [36, IV.6, p. 82], and so $H_{I,K}$ does not admit any nonconstant maps to an abelian variety.
The version of the modularity conjecture to be discussed here can be expressed in two forms. The weak form, which could still be considered algebraic, is that the Galois representation on $\operatorname{H}^{d}_{{\mathrm{\acute{e}t}}}(H_{I,K})$ has a subquotient isomorphic to $\otimes_{i=1}^{d}\operatorname{H}^{1}_{{\mathrm{\acute{e}t}}}\rho_{i}(E)$, where as above the $\rho_{i}$ are the embeddings of $K$ into ${\mathbb{R}}$. The strong form was first stated by Hamahata:
Conjecture 1.9.
[14, Section 0, p. 194] Let $E$ be an elliptic curve over a real quadratic field $K$ with conductor $I$. Then there is a correspondence between $H_{I,K}$ and $K_{E}=(E\times E^{\sigma})/\pm 1$ that induces an injection from the primitive part of $\mathcal{H}^{2}(K_{E})$ to that of $\mathcal{H}^{2}(H_{I,K})$.
Remark 1.10.
As Hamahata notes, this conjecture is inspired by the work of Oda [24]; we therefore refer to it as the Oda-Hamahata conjecture. It is an important special case of the Tate conjecture. An especially strong form of this conjecture would state that this correspondence can be defined over ${\mathbb{Q}}$, not just over $\bar{\mathbb{Q}}$.
Remark 1.11.
More generally, let $K$ be a totally real number field of arbitrary degree and let $E$ be an elliptic curve over $K$ of conductor $I$. Let $A=\prod_{i=1}^{[K:{\mathbb{Q}}]}E^{\sigma_{i}}$, where the $\sigma_{i}$ are the embeddings of $K$ into ${\mathbb{R}}$. Then $A$ admits an action of the group $G=({\mathbb{Z}}/2{\mathbb{Z}})^{d}$, where each generator acts as negation on one factor and the identity on the others. Let $G_{e}$ be the subgroup of elements whose sum is 00 (in particular, if $d=2$ we have $G_{e}=\{\pm 1\}$ and the quotient is the classical Kummer surface [7] as above). The quotient $A/G_{e}$ is a Calabi-Yau variety and the primitive part of $\operatorname{H}^{d}(A/G_{e})$ can be identified with $\otimes_{i}\operatorname{H}^{1}(E^{\sigma_{i}})$.. It is natural to generalize the Oda-Hamahata conjecture to this situation: namely, we ask for a correspondence between $A/G_{e}$ and $H_{I,K}$ inducing an injection from the primitive part of $\operatorname{H}^{d}(A/G_{e})$ to that of $\operatorname{H}^{d}(H_{I,K})$. For now this conjecture appears to be beyond computational attack for $d>2$.
Hamahata’s conjecture is known under some special circumstances: in particular, if $E$ is a ${\mathbb{Q}}$-curve, i.e., an elliptic curve isogenous to all of its Galois conjugates. This is not clearly stated or proved in the literature, but is known to experts. It is a consequence of the modularity of $E$, which is proved in [27, Corollary 6.2], subject to Serre’s conjecture on modular $2\times 2$ Galois representations, which is now a theorem of Khare and Wintenberger [17] More precisely, Ribet’s result is that $E$ is a quotient of $J_{1}(N)$; it follows that there is a dominant map $J_{1}(N)^{d}\to\prod_{i=1}^{d}E^{\sigma_{i}}$, where $d=[K:{\mathbb{Q}}]$. The domain of this map is related to $\operatorname{H}^{d}(X_{0}(N))$ by Hecke correspondences, and the equality of the levels follows from Serre’s formula for the level of the modular form, also proved by Khare and Wintenberger.
Our main goal in this paper is to prove some new cases of this conjecture. This advance is possible because of the recent release of software [2] for computing rings of Hilbert modular forms.
Remark 1.12.
Let ${\mathcal{O}}$ be an order in a totally real field and ${\mathcal{O}}^{+}$ the set of its totally nonnegative elements. Let ${\mathfrak{d}}_{\mathcal{O}}$ be the different of ${\mathcal{O}}$. A Hilbert modular form for ${\mathcal{O}}$ has a Fourier expansion in which the coefficients are indexed by elements of ${\mathcal{O}}^{+}\cap{\mathfrak{d}}_{\mathcal{O}}^{-1}$ and are invariant under multiplication by elements of ${\mathcal{O}}^{+}\cap{\mathcal{O}}^{*}$. In order to multiply two such expansions, one needs an algorithm for writing an element of ${\mathcal{O}}^{+}$ as a sum of two such elements in all possible ways. An efficient algorithm for doing so will be described in [1].
We state our main result here:
Theorem 1.13.
Let $K$ be a real quadratic field of narrow class number $1$ and $I$ an ideal of ${\mathcal{O}}_{K}$. Let $H_{I,K}$ be the associated Hilbert modular surface; assume that $I$ is an elliptic curve on the list in Theorem 1.2. Then there is an elliptic curve $E/K$ with conductor $I$ and an explicit correspondence between $H_{I,K}$ and $E\times E^{\sigma}$ (possibly defined only over a nontrivial extension of ${\mathbb{Q}}$) that identifies the transcendental parts of $H^{2}$ of the two surfaces.
Remark 1.14.
In particular, this theorem applies in all cases where $H_{I,K}$ is not of general type and $I\neq(1)$, except for $K={\mathbb{Q}}(\sqrt{13}),I=(2)$. In this case we do find a K3 surface that is in correspondence with the Hilbert modular variety (and in fact this has been known for some years, being [35, Theorem 9]) but it is quite difficult to establish an explicit correspondence with the desired Kummer surface. In two of the cases where $I=(1)$ the birational equivalence between the Hilbert modular surface and a K3 surface with canonical singularities is made explicit in [37, Appendix].
We have also considered a few cases in which the field $K$ has narrow class number $2$. One of these will be discussed in Section 5.4.
Our method is to study the Baily-Borel compactification $H_{I,K}$ by means of computer algebra and exhibit the desired correspondence to the Kummer surface by explicit maps. In some cases the Kummer surface of $E\times E^{\sigma}$ is itself a quotient of $H_{I,K}$; in other cases this does not appear to be true and instead we find an auxiliary K3 surface $S$ that admits maps of finite degree from both $H_{I,K}$ and the Kummer surface of $E\times E^{\sigma}$. The methods are similar to those of [37, Appendix], except that we do not necessarily start with a variety birational to a K3 surface. In addition, the setup of [37] requires information about the ring of Hilbert modular forms that is not available here: in particular, that it can be generated by Borcherds products. Against this, we do not find equations for the Hirzebruch-Zagier cycles [36, Def. VI.1.4] on the Hilbert modular surfaces.
The Hilbert modular surfaces are not only interesting for their relation to elliptic curves over real quadratic fields; they are worth studying in their own right. In particular, some such surfaces, or their quotients by Atkin-Lehner involutions, furnish examples of surfaces of geometric genus 00 and small $h^{1,1}$. It is interesting to relate these to known examples of surfaces with these properties, such as the famous Godeaux and Campedelli surfaces (see [25], for example). As suggested by Hamahata [14, p. 194], we would expect that the Hilbert modular threefolds (and their Atkin-Lehner quotients) also give examples of threefolds with unusual properties. For example, in light of the work of Grundman [12, 13], some such threefolds at levels where the geometric genus is 00 are varieties of general type whose Hodge diamond is that of a rational variety. Hilbert modular threefolds with geometric genus $1$ may also be interesting examples of varieties of general type with small Hodge numbers, although the classification problem, already very difficult for surfaces, is presumably hopeless in this situation.
In other cases, the surface or a quotient has geometric genus $1$ but is of general type. It is known [23, Theorem 1] that if a surface $S$ is of geometric genus $1$ then there is a K3 surface $K_{S}$ and a correspondence between $S$ and $K_{S}$ identifying the transcendental part of $H^{2}_{{\mathrm{\acute{e}t}}}$ for the two surfaces. However, finding such a K3 surface explicitly is a difficult and interesting problem in general. In some such cases the quotient by some or all of the Atkin-Lehner involutions is a surface with $p_{g}=1$ and $\kappa<2$ and we can find the desired correspondence, but otherwise there is no obvious method to use. This problem will be studied in future work.
Acknowledgments
This paper began when Wei Zhang asked John Voight whether advances in theory and computation in recent decades would make it possible to prove some new cases of Oda’s conjecture. Prof. Voight passed the question along to me; I thank him for doing so and for many interesting and informative discussions along the way, including some essential references. In addition, this paper could not have been written without the software [2]. I am grateful to its developers and especially to Eran Assaf and Edgar Costa for their explanations of how best to use the code. My understanding of modularity, and in particular of the distinction between the algebraic and geometric viewpoints, owes much to discussions with Jared Weinstein. I also thank Damián Gvirtz-Chen for suggesting an interesting calculation.
While writing this paper I benefited from the hospitality of ICERM and the support of the Simons Collaboration on Algebraic Geometry, Number Theory, and Computation (grant $546235$), as well as the liberality of the Tutte Institute for Mathematics and Computation in permitting me to take leave from my employment there and the opportunity to visit the Department of Pure Mathematics at the University of Waterloo.
2. General approach to constructing a K3 surface
We start by establishing the notation.
Notation 2.1.
Let $K$ be a real quadratic field, let ${\mathcal{O}}_{K}$ be its maximal order, and let $I$ be an ideal of ${\mathcal{O}}_{K}$. We define $M_{I}=M_{I,K}$ to be the ring of Hilbert modular forms of level $I$ and parallel weight. Further, we let $H_{I}=H_{I,K}=\operatorname{Proj}M_{I,K}$ be the well-known Baily-Borel compactification [36, II.7] of the Hilbert modular surface of level $I$.
In this paper we only use forms of parallel weight, so we will abuse notation by writing $w_{i}$ instead of $(w_{i},w_{i})$ for the weight of a Hilbert modular form. In addition, the presence of $-I_{2}$ in $\Gamma_{0}(N)$ implies that all forms are of even weight.
Definition 2.2.
Let $f_{1},\dots,f_{n}$ be a set of modular forms that generate $M_{I}$. Let their weights be $w_{1},\dots,w_{n}$. Then we define $\mathbb{P}_{I}=\mathbb{P}_{I,K}$ to be the projective space over ${\mathbb{Q}}$ with variable degrees $w_{i}/2$.
As above, $H_{I,K}$ is naturally a subvariety of $\mathbb{P}_{I}$. The ring $M_{I}$ is computed by the software package [2].
2.1. Atkin-Lehner involutions
The ring $M_{I}$ and the surface $H_{I}$ are rather complicated. When the level is not $1$, they always admit Atkin-Lehner involutions, and the quotient of $H_{I}$ by these, corresponding to the fixed subring of $M_{I}$, is simpler. Recall [11] that if $I=I_{1}I_{2}$ with $I_{1}+I_{2}=(1)$, then there is an Atkin-Lehner involution $w_{I_{1}}$ on $M_{I}$ and hence on $H_{I}$ corresponding to $I_{1}$. One can describe $w_{I}$ in terms of the moduli functor of abelian varieties with level structure represented by $H_{I}$. Namely, a point corresponds to an abelian surface with real multiplication by ${\mathcal{O}}_{K}$ and a subgroup isomorphic to ${\mathcal{O}}_{K}/I$ as an ${\mathcal{O}}_{K}$-module, and $w_{I_{1}}$ replaces the pair $(A,S)$ with $(A/I_{2}S,(S+A[I_{2}])/I_{2}S)$.
This abstract definition, however, is not very useful for computing the Atkin-Lehner involution in practice. Instead we use the description in terms of modular forms. The software computes the action on newforms. For oldforms, the following description, which follows from the $\operatorname{GL}(2)$ Oldforms Theorem of [28, p. 4], is sufficient for our purposes.
Proposition 2.3.
With notation as above, let $P$ be a prime ideal of ${\mathcal{O}}_{K}$, let $J$ be an ideal, and let $f$ be a Hecke eigenform of level $J$ and weight $w$. Consider the level-raising operator $a_{P^{i}}$ from forms of level $J$ to forms of level $JP^{d}$, where $0\leq i\leq d$. Let $w_{P}$ be the Atkin-Lehner operator associated to the highest power of $P$ dividing $JP^{d}$. Then $w_{P}(a_{P^{i}}f)=sp^{(d-2i)(w/2)}a_{P^{d-i}}f$, where $s=1$ if $P\nmid J$ and is equal to the Atkin-Lehner eigenvalue on $f$ otherwise.
Thus we may compute the Atkin-Lehner operator on the entire space of modular forms by expressing the basis given by the software in terms of Eisenstein series (for which again the action is easily described) and the images of newforms of various levels under degeneracy maps. Although the given basis does not consist of such forms, we may change to a different basis that does. There is no ambiguity because we can compute enough Hecke eigenvalues to determine the expression uniquely. We can verify that our expression is correct by checking that it gives an automorphism of $H_{I}$.
2.2. Projective embedding and cusps
In computing the Atkin-Lehner operators we determined the relation between the set of generators of $M_{I}$ used by [2] and bases of Eisenstein series and cusp forms. Thus we may identify the cusps of the surface as the points at which all cusp forms vanish. (The cusps belong to the singular locus of $H_{I}$, which in theory we could find by the standard Jacobian ideal computation, but this is infeasible in practice in all but a few cases.) This gives an additional test for our computation of Atkin-Lehner operators, since these should permute the cusps.
We now choose an integer $n$ and use it to embed $H_{I}$ into projective space by ${\mathcal{O}}(n)$. If we are considering the quotient by some or all of the Atkin-Lehner involutions, then rather than using all of ${\mathcal{O}}(n)$ we use only the subspace fixed by the involutions. In addition, we record the images of the cusps.
It is much easier to use computer algebra systems to work with a surface if the singularities of the surface are canonical; in other words, if they are of ADE type, or Du Val singularities [26, (1.2)]. However, the cusp singularities never have this property, and not all elliptic points [36, p. 15] do. Thus we would like to resolve these singularities.
Zariski showed that the singularities of a surface can always be resolved by alternating between normalizing and blowing up isolated singularities. On a computer, both of these are rather unpleasant procedures. As a substitute for blowing up, we use the projection. That is, let $S\subset\mathbb{P}^{n}$ be a surface with an isolated singular point $P$. Projection away from $P$ gives a rational map $S\dashrightarrow\mathbb{P}^{n-1}$, whose image we will denote by $S_{P}$. Abstractly we think of the situation as follows. Let $S^{\prime}$ be the partial resolution of $S$ at $P$ given by blowing up once, let $H$ be the hyperplane class, and let $E$ be the exceptional divisor. Then $S_{P}$ may be viewed as the embedding of $S^{\prime}$ by means of the divisor class $H-E$.
We know $E$ abstractly by the ideas of Hirzebruch [36, Chapter II], and it is computed by [2] as a cycle of rational curves of known self-intersection, each one intersecting the adjacent ones. In particular, this can be used to predict the degree of $S_{P}$. This is valuable, because a K3 surface embedded into $\mathbb{P}^{n}$ by a complete linear system always has degree $22n-$ 2, and so a projection that reduces the degree by $3$ or more means that we are making progress (unless the projection itself is not a birational equivalence).
In some cases, after repeated projections away from the singularities, we obtain a surface $S$ that is singular along a line $L$. A variety that is singular in codimension $1$ is not normal [18, I.4], so naturally we want to project away from $L$. The image $S_{L}$ of $S$ under projection away from $L$ is in a space of dimension $2$ less and satisfies $\deg S-\deg S_{L}\geq 5$, so such an operation always brings us closer to the degree for a standardly embedded K3 surface.
It can happen that we run out of room, in the sense that projection away from a singular point or line gives a map of degree greater than $1$. This is undesirable, because the image may no longer have the desired cohomology. When it happens, we apply the $2$-fold Veronese embedding and proceed with our projections as before.
Even after we have a K3 surface, we may want to make further birational transformations in order to improve the model. In particular, the lower the degree of the singularities of a K3 surface, the easier it is to work with its linear systems; the tradeoff is that it is more difficult in a higher-dimensional projective space. As before, the basic idea is to project away from the worst singularities and use the Veronese embedding when necessary to find more room. Typically it is fairly easy to compute on a model in $\mathbb{P}^{7}$ with only $A_{1},A_{2}$ singularities, or in $\mathbb{P}^{5}$ when there are only $A_{1},A_{2},A_{3}$ points.
2.3. Tricks for finding images of projections and singular points
In principle the problem of finding the image of a rational map of varieties can be solved algorithmically by means of an algorithm using Gröbner bases [8, Sections 3.1–3.3]. In practice this algorithm, as implemented in Magma [5], may require more computer time or memory than is available, and so we may need to find the image in other ways.
First, even if the map is defined by linear equations, as is the case for a projection map, the implementation does not appear to take advantage of this but simply treats it as a general map defined by arbitrary polynomials. However, in this case, the image can be found by calculating an elimination ideal of the ideal defining the source, rather than of the graph as in the general method. To state the result concretely:
Proposition 2.4.
Let $V\subseteq\mathbb{P}^{n}$ be the subvariety defined by the ideal $I\subseteq K[x_{0},\dots,x_{n}]$. Fix $k\leq n$ and let $\pi$ be the rational map $\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{k}$ defined by $(x_{0}:\dots:x_{k})$. Then the Zariski closure of $\pi(V)$ is defined by the elimination ideal $K[x_{0},\dots,x_{k}]\cap I$.
This is almost the definition of the image, but it is a very useful fact since it is much faster to compute an elimination ideal of $I$ than of the graph of $\pi$, as is required by the general algorithm. (Note that this method can be used even when the map is defined by linear equations that are not monomials, by composing with a linear automorphism of $\mathbb{P}^{n}$.) However, in some cases even this method fails in practice, or we need to find the image of a map defined by polynomials of degree greater than $1$. Then we are forced to resort to another method, which is based on an even simpler observation:
Lemma 2.5.
Let $V$ be a variety and $\pi:V\dashrightarrow W$ a rational map. Let $C\subseteq V$ be a Zariski closed subset. Then $\pi(C)\subseteq\pi(V)$.
This trivial statement is useful because it is generally much faster to compute the image of a 00-dimensional scheme. Let us state another proposition that shows that this idea can be powerful.
Proposition 2.6.
Let $V$ be a variety and $C_{0},C_{1},\dots$ an infinite sequence of Zariski closed subsets of $V$, the Zariski closure of whose union is $V$. Let $\pi:V\dashrightarrow W$ be a dominant rational map, and let $W_{k}=\cup_{i=1}^{k}\pi(C_{i})$. Then:
  1. (1)
    For all $k$, every polynomial vanishing on $W$ vanishes on $W_{k}$.
  2. (2)
    For all $d>0$, there exists $k_{d}$ such that every polynomial of degree $d$ vanishing on $W_{k_{d}}$ vanishes on $W$.
Proof.
The first statement holds because $W_{k}\subseteq W$ by Lemma 2.5. For the second, let $D_{k,d}$ be the dimension of the space of homogeneous polynomials of degree $d$ vanishing on $W_{k}$ modulo those in $I(W)$. Clearly $D_{k,d}\geq D_{k+1,d}\geq 0$ for all $d,k$. We show that if $D_{k,d}\geq 0$ then $D_{k+m,d}<D_{k,d}$ for some $m$, from which the result follows. Indeed, let $p$ be a homogeneous polynomial of degree $d$ in $I(W_{k})\setminus I(W)$. By hypothesis, the Zariski closure of $\cup_{i\in{\mathbb{N}}}\pi(C_{i})$ is $W$, so it is not contained in $W\cap(p=0)$. Thus for some $m$ we must have $p\notin I(\pi(C_{k+m}))$ and hence $p\notin I(W_{k+m})$ and $p\notin D_{k+m,d}$. ∎
This suggests the following empirical approach to finding the image of a map $\pi:V\dashrightarrow\mathbb{P}^{n}$:
Algorithm 2.7.
  1. (1)
    To start, let $I$ be the empty subscheme of $\mathbb{P}^{n}$.
  2. (2)
    Choose a linear subspace $L$ of the ambient space of $V$ of codimension $\dim V$ and find $\pi(L\cap V)$.
  3. (3)
    Replace $I$ with $\pi(L\cap V)\cup I$ and observe the degrees of the generators of $I$.
  4. (4)
    If it appears that the low-degree generators have stabilized, then guess that $\pi(V)$ is defined by them. If not, return to Step 2.
Of course, this method is incapable of proving that $\pi(V)$ is as we believe it to be. But in many cases, once we have a candidate $C$ for $\pi(V)$, we can verify that it contains $\pi(V)$ (this just means that every generator of the ideal of our candidate pulls back to an element of $I(V)$). Often further information is available that allows us to prove that our candidate $C$ really is equal to $\pi(V)$: for example, the restriction of $\pi$ to a map $V\to C$ is invertible, or $C$ is irreducible and $\pi(V)$ contains two distinct codimension-$1$ subvarieties of $C$.
We now discuss how to find the singular points on a projective variety $V$ defined over ${\mathbb{Q}}$. Again, this can be done in principle by means of a Jacobian ideal computation, but when the variety is defined by a large number of equations in a projective space of high dimension that is impractical.
To illustrate the method in the simplest case, let us suppose that $V$ has exactly one singular point. For almost all primes $p$, the reduction of $V$ mod $p$ will have the same property, and we can find the point mod $p$ by enumerating the ${\mathbb{F}}_{p}$-points of the reduction and checking them for singularity (deciding whether a point is singular is a simple computation if we know the dimension of $V$). By looking for rational numbers that reduce mod the various $p$ to the coordinates of these points, we guess the coordinates of the point over ${\mathbb{Q}}$. Again, this is not a proof, but we may easily check whether our candidate is actually on $V$ and, if so, whether it is singular.
Remark 2.8.
There is a well-known method for finding $a\in{\mathbb{Q}}$ reducing to $a_{i}\mod p_{i}$ for primes $p_{1},\dots,p_{n}$ and $a_{i}\in{\mathbb{F}}_{p_{i}}$ for $1\leq i\leq n$. To wit, let $L$ be the lattice ${\mathbb{Z}}^{2}$ and for each $i$ consider the sublattice $L_{i}$ generated by $pL$ and $(a_{i},1)$. Let $M=\cap_{i=1}^{n}L_{i}$. If $a=b/c$ in lowest terms then $(b,c)\in M$, and if $n$ is large enough then it will almost certainly be the shortest vector. Reduction of $2$-dimensional lattices is straightforward, so we can find $a$.
More generally, there might be more than one singular point on $V$ and they might not be rational. This may cause problems because we would not know which collections of reductions come from the same singular point; even if there are only $2$ singular points trying all the possibilities leads to a combinatorial explosion. In some cases the points will have some properties that make it clear which reductions go together, such as some conspicuous 00 coordinates, or we will already know some singular points. Otherwise we may have to consider the reduced 00-dimensional schemes supported at the singular points mod $p$ and interpolate the coefficients in the Gröbner basis of the ideals defining them.
If the points are not rational, then of course for some primes we will only be able to find their reductions by passing to an extension of ${\mathbb{F}}_{p}$, which is undesirable. Fortunately in this work it was rarely necessary to find singular points not defined over a quadratic field. This case can be recognized by the singular points appearing on the reductions modulo half of the primes. There are a few more complicated cases with singular points defined over different fields; however, the quadratic fields of definition are easily recognized and so one can guess a formula for the number of singular points over ${\mathbb{F}}_{p}$ in terms of quadratic residue symbols. By considering the reductions modulo primes for which exactly one of the symbols is $1$ one can find the points.
3. Verification of modularity
Let us now suppose that we have found two K3 surfaces: a quotient $K_{0}$ of the Hilbert modular surface $H_{I}$ and the Kummer surface $K_{1}$ of $E\times E^{\sigma}$. We would like to show that these are in correspondence, which in practical terms means that there is a sequence of K3 surfaces $K_{0}=L_{0},L_{1},\dots,L_{n}=K_{1}$ and, for all $0<i<n$, a rational map of finite degree $L_{i}\to L_{i+1}$ or $L_{i+1}\to L_{i}$. Essentially we know of four ways to construct rational maps, three of which are based on genus-$1$ fibrations.
Let $S$ be a K3 surface defined over ${\mathbb{Q}}$, and let $\pi:S\to\mathbb{P}^{1}$ be a map whose general fibre is a smooth curve of genus $1$.
Notation 3.1.
Let $L$ be a lattice. For $r\in{\mathbb{Q}}^{+}$ let $rL$ be the sublattice of $L\otimes{\mathbb{Q}}$ whose vectors are the $rv$ for $v\in L$, and let $L(r)$ be the lattice whose underlying abelian group is that of $L$ but with the pairing $(x,y)_{L(r)}=r(x,y)_{L}$. (To clarify this notation, we point out that $rL$ and $L(r^{2})$ are isometric.)
  1. (1)
    Let $p$ be a prime and suppose that the $p$-torsion of the general fibre of $\pi$ has a nontrivial subgroup $T$ of order $p$ defined over ${\mathbb{Q}}(t)$. In this case we may define an isogeny on the general fibre of $\pi$ by quotienting by $T$; this spreads out to a map from $S$ to another K3 surface. The effect of this on the transcendental lattice $T_{S}=\operatorname{H}^{2}(S)/\operatorname{Pic}S_{\bar{\mathbb{Q}}}$ of $S$ is to replace it by a lattice between $T_{S}(1/p)$ and $T_{S}(p)$. (This exact statement is not easy to find in the literature. It follows from remarks in [4, Section 2.4] that a map $X\dashrightarrow Y$ of K3 surfaces of degree $n$ induces an embedding $T_{Y}(n)\hookrightarrow T_{X}$, whence the given statement follows from the existence of a dual isogeny.)
  2. (2)
    If $\pi$ has no sections, then let $d$ be the smallest degree of a multisection (which is the GCD of the degrees of intersections of curves on $S$ with a fibre of $\pi$). The map from the generic fibre to its Jacobian spreads out to a map of degree $d^{2}$ from $S$ to another K3 surface $\operatorname{Jac}_{\pi}(S)$. If $E$ is the class of a fibre of $\pi$, then we have $\operatorname{Pic}\operatorname{Jac}_{\pi}(S)=\operatorname{Pic}S[E/d]$ [16, Lemma 2.1], at least after base change to $\bar{\mathbb{Q}}$ (in other words, this is not an isomorphism of Galois module structures).
  3. (3)
    It can happen that $\pi$ has no sections over ${\mathbb{Q}}$ but does over a larger number field $K$. In this case $\operatorname{Jac}_{\pi}(S)$ is $K$-isomorphic to $S$ and so we have $\operatorname{Pic}S_{/K}\cong\operatorname{Pic}\operatorname{Jac}_{\pi}(S)_{/K}$. However, the submodule of $\operatorname{Pic}\operatorname{Jac}_{\pi}(S)$ fixed by $\operatorname{Gal}({\mathbb{Q}})$ is larger than for $S$. This can be a useful operation because it enlarges the range of fibrations available to us, thus creating new maps of types (1), (2).
  4. (4)
    If $S$ is a quotient of an abelian surface $A$ by a finite group $G$ of automorphisms and $\phi:A\to A^{\prime}$ is an isogeny whose kernel is $G$-invariant, then $\phi$ descends to a rational map $S\to S^{\prime}$. In particular this applies if $S$ is a Kummer surface, the condition on $\ker\phi$ being automatic in that case. We can understand the relation between the transcendental lattices of $S$ and $S^{\prime}$ induced by such a map from that between $T_{A},T_{A^{\prime}}$, since $T_{S}=T_{A}(2)$ and $T_{S^{\prime}}=T_{A^{\prime}}(2)$.
Remark 3.2.
Note that in our situation the geometric Picard rank of $K_{1}$ is at least $18$, so there is always an abundance of genus-$1$ fibrations.
Remark 3.3.
The geometric Picard rank of a K3 surface is invariant under finite maps [4, Theorem 1.1]. There are additional invariants in the cases of rank $18$ and $19$ that we are concerned with here. Namely, when the rank is $18$ (or any even integer), then the class of the discriminant of the Picard lattice in ${\mathbb{Q}}^{*}/({\mathbb{Q}}^{*})^{2}$ is invariant under all three types of maps described above. When the rank is $19$, the intersection form on $T_{S}$ defines a conic in $\mathbb{P}^{2}$, whose ${\mathbb{Q}}$-isomorphism class is invariant under the three types of maps.
Remark 3.4.
Let $K_{0},K_{1}$ be K3 surfaces for which we would like to find a correspondence. Let the $D_{i}$ be the discriminants of the respective geometric Picard lattices, and let $P$ be the set of primes such that $v_{p}(D_{0}/D_{1})$ is odd. If the rank is even and $P$ is nonempty, the task is hopeless; if the rank is odd, we start by finding a $p$-isogeny for each prime dividing $p$ (if this is not possible, the task again seems hopeless). Thus we assume that $P$ is empty.
In general it is easier to work with K3 surfaces that have a large part of their Picard lattice defined over ${\mathbb{Q}}$; this makes it easier to find elliptic fibrations and maps as defined above. This generally goes with a smaller Picard discriminant. Thus we want to apply maps that reduce the discriminant if possible; for isogenies this can be difficult to predict, but it always happens when we take a Jacobian. On the other hand, if we cannot find any fibrations without a section defined over ${\mathbb{Q}}$, it may still be possible to describe a fibration with a section but none over ${\mathbb{Q}}$ and apply a step of type 3, which then allows us to continue by a step of type 1 or 2.
Another advantage of reducing the discriminant is that there are only a few lattices of small discriminant, and each one supports only a small number of frame lattices for elliptic fibrations [29, (8.6)]. This makes it easier to recognize the isomorphism of K3 surfaces.
Remark 3.5.
Let us suppose that we have found a Weierstrass equation for the general fibre of an elliptic fibration. It may be defined by polynomials with unreasonably large coefficients, causing future computations to be slow. This situation can often be remedied as follows: find a minimal integral model and let $D(t)$ be the discriminant, ignoring any contribution at infinity so that this is a polynomial. Find an automorphism of $\mathbb{P}^{1}$ under which the numerator of the pullback of $D$ has smaller coefficients. (In Magma, this amounts to minimizing and reducing the hyperelliptic curve defined by $y^{2}=D(t)$.) Changing coordinates on the base by means of the same linear transformation often results in a much more usable equation, and twisting by a large square common factor of the coefficients is a further improvement.
4. Examples
4.1. Summary
In this paper we are mainly concerned with the cases where $p_{g}(H_{I,K})=1$. Up to Galois conjugacy, Hamahata [14] lists $14,6,4,2$ such surfaces over the real quadratic fields of discriminant $5,8,13,17$ respectively. (Such surfaces also exist over fields whose strict class number is not $1$, and over the fields of discriminant $29,37,41$ at level $1$, but these are not his concern.) We summarize his results on the Kodaira dimensions in the following table:
$\kappa=0$ $\kappa=1$ $\kappa=2$ $\kappa\in\{1,2\}$
$D=5$ 0 0 9 5
$D=8$ 1 2 2 1
$D=13$ 1 1 2 0
$D=17$ 2 0 0 0
Remark 4.1.
In this table we have corrected an error in [14]: as we will see in Section 4.7, the surface $H_{{\mathfrak{p}}_{17},{\mathbb{Q}}(\sqrt{2})}$ has Kodaira dimension $1$. This means that the argument in [14, 6.9] is invalid; in fact Hamahata’s $d_{m}(4;1,1)$ is approximately $m^{2}/2$, so that the lower bound for $P_{m}$ is $O(m)$. I thank Eran Assaf for explaining this to me.
Remark 4.2.
The cases of ambiguous Kodaira dimension from [14] can all be resolved. In particular, let us define the standard model of a Hilbert modular surface (a standard concept, but not a standard term) to be that obtained by resolving all singularities and then recursively contracting all Hirzebruch-Zagier cycles and components of resolutions of self-intersection $-1$. If the Kodaira dimension is nonnegative, then no model has intersecting $-1$-curves, so this is unambiguous. Such a model is believed to be minimal, but this is not known in general.
The software [2] can compute the self-intersection of the canonical divisor on the standard model. In five of the six cases left undecided by Hamahata, namely $(6),3{\mathfrak{p}}_{5},{\mathfrak{p}}_{5}{\mathfrak{p}}_{11},4{\mathfrak{p}}_{5}$ over ${\mathbb{Q}}(\sqrt{5})$ and $2{\mathfrak{p}}_{7}$ over ${\mathbb{Q}}(\sqrt{2})$, it is positive, so the Kodaira dimension cannot be 00 or $1$ and must be $2$ since $\operatorname{h}^{2,0}\neq 0$. On the other hand, the self-intersection is 00 for ${\mathfrak{p}}_{31}$ over ${\mathbb{Q}}(\sqrt{5})$. Again, this does not show that the surface is not of general type, because the model is not known to be minimal. We will discuss this example in detail in Section 4.8, showing that the Kodaira dimension is $1$.
Remark 4.3.
The surfaces of general type are extremely interesting, but they seem to be very difficult for the methods of this paper. As mentioned in the introduction, a general theorem of Morrison [23] asserts the existence of a correspondence between a surface with $\operatorname{h}^{2,0}=1$ and a K3 surface. However, except in some special cases such as Todorov surfaces (double covers of K3 surfaces that were first studied as examples of surfaces of general type that do not satisfy a Torelli theorem), the correspondence is very hard to find. Some of these surfaces admit Atkin-Lehner quotients that are not of general type; instead they are elliptic surfaces of a very interesting kind. This applies to $H_{{\mathfrak{p}}_{19},{\mathbb{Q}}(\sqrt{5})}$, for example.
Hamahata also studies the surfaces with $p_{g}=0$. Since these are not associated with any modular forms they are not relevant for the Oda-Hamahata conjecture, but many of them are of general type and are thus relevant to the important tradition of constructing surfaces of general type with $p_{a}=p_{g}=0$ and small $K^{2}$. We intend to study these two situations in more detail in future work.
Of the $7$ cases with $p_{g}=1$ where the surface was known not to be of general type, there are only $2$ where the level is not Galois-stable and thus the conjecture is not already known. As mentioned above, the surface $H_{{\mathfrak{p}}_{31},{\mathbb{Q}}(\sqrt{5})}$ has Kodaira dimension $1$, and we will show that $H_{2{\mathfrak{p}}_{7},{\mathbb{Q}}(\sqrt{2})}$, another surface left undecided by Hamahata, is of general type. These and other examples will be discussed in Section 5.
Remark 4.4.
There are some cases in which $H_{I,K}$ has $p_{g}=0$ but is nevertheless related to a K3 surface. For example, if $I=(3)$ and $K={\mathbb{Q}}(\sqrt{5})$, then $H_{I,K}$ is an Enriques surface and its unramified double cover is a K3 surface of Picard rank $20$ and discriminant $-240$. Thus it is related to the Kummer surface of $E\times E^{\sigma}$, where $E$ is an elliptic curve with complex multiplication by $\sqrt{-15}$, in particular a curve in the class 2.2.5.1/81.1-a with conductor $(9)$.
Definition 4.5.
Let ${\mathcal{I}}$ be an isogeny class of elliptic curves (over a number field, or otherwise constrained to be finite). If $E_{1},E_{2}\in{\mathcal{I}}$, let $d(E_{1},E_{2})$ be the minimal degree of an isogeny from $E_{1}$ to $E_{2}$. (Note that the distance function $\log d$ makes ${\mathcal{I}}$ into a metric space.) The diameter $\operatorname{dia}({\mathcal{I}})$ is defined to be $\max_{E_{1},E_{2}\in{\mathcal{I}}}d(E_{1},E_{2})$.
The Picard group of the K3 surfaces that are quotients of the Hilbert modular surface may be of interest. Empirically the following statement seems to hold:
Claim 4.6.
Let $H_{I,K}$ be a Hilbert modular surface and let $S$ be a K3 surface which is a quotient of $H_{I,K}$ not factoring through any other map to a K3 surface and associated to an isogeny class ${\mathcal{I}}$. The Picard discriminant of $S$ is a very smooth number (usually a small power of $2$) times $\operatorname{dia}({\mathcal{I}})$.
In particular, if $\operatorname{dia}({\mathcal{I}})$ is divisible by a prime greater than $5$, we expect it to be very difficult to find a correspondence between $S$ and a Kummer surface.
We now proceed to a detailed examination of some individual examples. The code that supports the claims made in the examples of this and the next section is available at [21].
4.2. Level ${\mathfrak{p}}_{2}^{5}$ over ${\mathbb{Q}}(\sqrt{2})$
Our first example is very much simplified by the fact that the associated elliptic curve is not merely a ${\mathbb{Q}}$-curve but in fact the base change of an elliptic curve over ${\mathbb{Q}}$ with complex multiplication. In particular, the elliptic curve $y^{2}=x^{3}-x$ has conductor ${\mathfrak{p}}_{2}^{5}$ over ${\mathbb{Q}}(\sqrt{2})$, with the reduction at ${\mathfrak{p}}_{2}$ being of type $I_{3}^{*}$. However, let us begin with the Hilbert modular surface. We calculate that the ring of Hilbert modular forms is generated by $9$ forms of weight $2$ and $1$ of weight $4$. The number $9$ comes from the $8$ cusps giving $8$ Eisenstein series and the $1$ cusp form corresponding to the elliptic curve.
According to [14, 6.11] the Hilbert modular surface $H_{{\mathfrak{p}}_{2}^{5},{\mathbb{Q}}(\sqrt{2})}$ is an honestly elliptic surface (i.e., not a K3 or rational elliptic surface). We will go directly to the quotient by the Atkin-Lehner involution. Since the space of weight-$4$ cusp forms with Atkin-Lehner eigenvalue $-1$ is contained in the space of products of weight-$2$ modular forms, we need a form in the $+1$ eigenspace, meaning that the quotient is naturally in weighted projective space $\mathbb{P}(1^{5},2)$. The quotient surface $S$ has $4$ cusps.
Proposition 4.7.
The image of $S$ by the map given by forms of degree $2$ vanishing on the cusps modulo those in the ideal of $S$ is a surface $S_{9}$ of degree $24$ in $\mathbb{P}^{9}$ in which the cusps resolve to singular lines.
Proof.
This is calculated by means of Algorithm 2.7. ∎
Projection away from any one of these lines gives a surface of degree $18$ in $\mathbb{P}^{7}$; since they are disjoint, one heuristically expects projection from the $\mathbb{P}^{7}$ that they span to give a “surface of degree 00 in $\mathbb{P}^{1}$”. This can be interpreted as a K3 surface with a genus-$1$ fibration, which prompts the next proposition.
Proposition 4.8.
The map $S_{9}\to\mathbb{P}^{1}$ given by the linear forms vanishing on all four singular lines has fibres of degree $16$ and arithmetic and geometric genus $1$.
Proof.
Once guessed this is easily checked in Magma. See [21]. ∎
By projecting the generic fibre away from the union of three of the four degree-$4$ divisors obtained by intersecting it with lines we find a quartic model in $\mathbb{P}^{2}$ singular at two points $P_{1},P_{2}$. The linear system of quadrics vanishing at $P_{1},P_{2}$ gives a map to $\mathbb{P}^{3}$ whose image is a smooth intersection of two quadrics in $\mathbb{P}^{3}$ by the map given by quadrics in $\mathbb{P}^{2}$ vanishing on the singular points. Taking the Jacobian of the generic fibre we define an elliptic curve $\mathcal{E}$ over ${\mathbb{Q}}(t)$ that spreads out to a K3 surface.
Remark 4.9.
We do not know whether the generic fibre admits any ¯$\bar{\mathbb{Q}}(t)$-rational divisors of degree between 00 and $4$, but we suspect not.
In fact, this elliptic curve has a simple description. Up to changes of coordinates on the base, it is the unique elliptic curve over $\mathbb{P}^{1}$ with fibres of types ${I_{1}^{*}}^{2},I_{8},I_{2}$; the torsion subgroup has order $4$, the rank must be 00 since the known curves already account for Picard rank $20$, and so the Picard discriminant is $-16$. Applying the method of Remark 3.5 we compute the equation
$y^{2}=x^{3}+(-2t^{3}+12t^{2}-2t)x^{2}+(t^{6}+4t^{5}+6t^{4}+4t^{3}+t^{2})x.$
Over $\bar{\mathbb{Q}}$ the general theory of K3 surfaces of rank 20 [31] can easily be used to show that such a surface is in correspondence with $\mathcal{K}(E\times E)$, where $E:y^{2}=x^{3}-x$ is an elliptic curve with complex multiplication by ${\mathbb{Z}}[i]$. In particular, one checks that the transcendental lattice is isomorphic to that of the lattice ${\langle 4\rangle}^{2}$ generated by two orthogonal vectors of norm $4$, and the result follows by [32, Section 3]. However, over ${\mathbb{Q}}$ our work is not quite done.
We remark that all components of reducible fibres are rational. Indeed, the rational $4$-torsion means that every component of a reducible fibre that meets a multiple of the $4$-torsion section must be rational, and none of the fibres admits any nontrivial permutation of its components preserving intersections and fixing all of these curves. On the other hand, the points of intersection of the components of the $I_{2}$ are not rational, being instead defined over ${\mathbb{Q}}(i)$.
Counting points mod $p$ on $\mathcal{E}$, we find results consistent with the existence of a correspondence to $\mathcal{K}(E\times E)$, where as above $E$ is defined by $y^{2}=x^{3}-x$.
Proposition 4.10.
There is a fibration on $\mathcal{K}(E\times E)$ whose general fibre is isomorphic to the elliptic curve $E^{\prime}$ defined by $y^{2}=x^{3}+(t^{3}+t)x^{2}-4t^{4}x+(-4t^{7}-4t^{5})$.
Proof.
We write down a model of the Kummer surface in $\mathbb{P}^{4}$ and use it to determine a fibration with two fibres each of type $I_{2}^{*}$ and $I_{4}$. Concretely, our model is calculated by mapping $E\times E\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$ to $\mathbb{P}^{4}$ by sections of ${\mathcal{O}}(1,1)$ invariant under the product of the two negation maps. It has an $A_{3}$ singularity which is the image of $E\times 0\cup\{(0,0)\}\cup 0\times E$ and nine $A_{1}$ singularities under the points $(T_{1},T_{2})$, where $T_{1},T_{2}$ are points of $E$ of exact order $2$.
The fibration we seek has fibres of class $H-E_{1}-E_{2}-E_{3}-L_{1}-L_{2}$, where the $E_{i}$ are the components of the resolution of the $A_{3}$ and $L_{1},L_{2}$ are the images of $(O,(0:0:1)),((0:0:1),O)\in E\times E$ (more precisely, of the curves above these points when the indeterminacy of the rational map $E\times E\dashrightarrow\mathbb{P}^{4}$ is resolved). It is a straightforward calculation in Magma to determine the general fibre of this map and its Jacobian. ∎
At this point, it only remains to verify that, up to change of coordinates on $\mathbb{P}^{1}$, the elliptic curves $E,E^{\prime}$ are isomorphic. It is easy to find the correct change of coordinates, because the bad fibres must be in the same places for the curves to be isomorphic.
4.3. Level $(3)$ over ${\mathbb{Q}}(\sqrt{13})$
We now present an example that, although still relatively simple, is more complicated than that of Section 4.2 and illustrates many of the methods of calculation used in this paper. Namely, we consider the isogeny class 2.2.13.1/9.1-a of elliptic curves of conductor $(3)$ over ${\mathbb{Q}}(\sqrt{13})$. Again, in this case the conductor is Galois-stable and so the Oda-Hamahata conjecture is already known. Nevertheless, the explicit calculation is of some interest.
Throughout this section, let $K={\mathbb{Q}}(\sqrt{13})$ and let $I=(3)$. According to Hamahata [14], the Hilbert modular surface $H_{I,K}$ has Kodaira dimension $1$ and hence has a unique genus-$1$ fibration. It would be possible to find this fibration; the Jacobian would be the desired K3 surface. Instead we begin by determining the Atkin-Lehner involutions. Since $3$ splits in $K$, there are two.
Proposition 4.11.
The subspace of ${\mathcal{O}}(3)$ (i.e., the space of modular forms of weight $6$) fixed by both Atkin-Lehner involutions modulo forms vanishing on the Hilbert modular surface has dimension $11$. A basis defines a map to $\mathbb{P}^{10}$ whose image is a surface $S_{10}$ of degree $24$. The four cusps of $H_{I,K}$ are permuted transitively by the Atkin-Lehner involutions and hence map to a single singular point of $S_{10}$.
Proof.
This is computed in Magma with the help of Algorithm 2.7 as has already been explained. For the last statement, we note that since the level is squarefree the cusps are in natural bijection with the divisors of $N$, and the Atkin-Lehner operator $w_{I}$ takes the cusp $c_{J}$ to $w_{IJ}$ if $IJ|N$ or $w_{IJ/N}$ if not, just as in the classical setting. ∎
We find the cusps on $H_{I,K}$ as the points where all cusp forms vanish. The cusp on $S_{10}$ is resolved by a cycle of rational curves of degrees $-5,-2,-2$ [2]. Let these curves be $E_{1},E_{2},E_{3}$. Then $({\mathcal{O}}(3)-(E_{1}+E_{2}+E_{3}))^{2}=24+(-2-2-5+3\cdot 2)=21$, since $({\mathcal{O}}(3),E_{i})=0$ for $i=1,2,3$. Thus, when we project away from the cusp, we obtain a surface $S_{9}$ of degree $21$.
Proposition 4.12.
The surface $S_{9}$ has at least two singular points, including the point $E_{9}$ which is the image of the elliptic points of $H_{I,K}$ by the map to $S_{10}$ followed by projection away from the cusp and the point $C_{9}$ which is in the image of the cusp.
Proposition 4.13.
Let $S_{8}$ be the projection of $S_{9}$ away from $C_{9}$ and let $C_{8}$ be the image of $C_{9}$ in $S_{8}$. For $i=8,7,6$, inductively define $S_{i-1}$ to be the projection of $S_{i}$ away from $C_{i}$ and $C_{i-1}$ to be the unique singular point in the image of $C_{i}$. Then $S_{i-1}$ is a surface in $\mathbb{P}^{i-1}$ of degree $12i-$ 1.
Proposition 4.14.
$S_{5}$ is the blowup of a K3 surface in a single point.
Proof.
Using built-in Magma functions for computing canonical models we find a map to a K3 surface of degree $10$ in $\mathbb{P}^{6}$ whose inverse fails to be defined at a single point. The inverse image of that point can easily be checked to be a smooth curve of self-intersection $-1$. ∎
This K3 surface $S_{10}$ of degree $10$ in $\mathbb{P}^{6}$ is defined by $6$ quadrics as expected. By good fortune all of its singularities are of type $A_{1}$; there are two orbits over ${\mathbb{Q}}(\zeta_{12})$, one over ${\mathbb{Q}}(\sqrt{13})$, and two over ${\mathbb{Q}}$, making $12$ singular points in all.
In order to study $S_{10}$ more closely, we find some curves on it. One way to do this is to note that one of the ${\mathbb{Q}}$-rational singular points has the property that a line through it and one of the two singular points defined and conjugate over ${\mathbb{Q}}(\sqrt{13})$ lies on the surface. Another is to use the genus-$1$ fibrations defined by hyperplanes passing through one orbit of ${\mathbb{Q}}(\zeta_{12})$-rational singular points and one rational singular point. These are all fibrations with three $I_{3}$ fibres and two of type $I_{4}$, and so each one gives us $17$ rational curves contained in fibres. However, some of these are the curves in the resolutions of the $A_{1}$ singularities, and the overlap means that we have only found $12$ curves in total.
This is enough for now. We have a sublattice of the Picard group with $18$ generators (later we will find a $19$th on a related K3 surface).
Proposition 4.15.
There is a genus-$1$ fibration on $S_{10}$ defined over ${\mathbb{Q}}$ whose Jacobian is defined by
$y^{2}$ $=x^{3}+(-8t^{4}-40t^{3}-2t^{2}+648t+150)x^{2}+(16t^{8}+160t^{7}+$
$\qquad 408t^{6}-2680t^{5}-15607t^{4}-18240t^{3}+53626t^{2}-31400t+5625)x.$
Proof.
Using our knowledge of the Picard group of $S_{10}$ and the action of Galois on it, we identify a certain fibration with fibres of degree $4$. The individual reducible fibres of this fibration are only defined over ${\mathbb{Q}}(i)$, so we start by writing down equations for the fibration over ${\mathbb{Q}}(i)$ and then move to ${\mathbb{Q}}$ by a standard Galois descent.
Since the fibres are of degree $4$, Magma has no trouble determining the Weierstrass model of the Jacobian. Using the method of Remark 3.5, we compute the model above. ∎
This is an elliptic surface $\mathcal{E}$ with two singular fibres of type $I_{0}^{*}$ and one each of types $I_{6},I_{3},I_{2},I_{1}$. (It is therefore a quadratic twist of the second family on [15, p. 336], which is the universal elliptic curve with a $6$-torsion point.) Over ${\mathbb{Q}}(\sqrt{13})$ there is a section with $x$-coordinate $(13/9)\cdot(t^{2}+8t+25)\cdot(2t-1)^{2}$; its height is $2/3$, since it passes through nonzero components of the two $I_{0}^{*}$ fibres and a component of the $I_{6}$ at distance $2$ from the zero section (cf. [9, Table 1.19, Theorem 1.33]). We calculate that the torsion subgroup has order $2$, so the determinant formula [29, Corollary 6.39] shows that the discriminant of the Picard lattice is $964^{2}\cdot 6\cdot 3\cdot 2\cdot 2/3\cdot 1/4=$ 96. This is not exactly what we are looking for, since our isogeny class contains elliptic curves $3$-isogenous to their Galois conjugates.
Since we will need genus-$1$ fibrations on the K3 surface, we must now find a good projective model of $\mathcal{E}$. Applying the techniques of Section 2.2 we rapidly come to a model $S_{3}$ in $\mathbb{P}^{3}$ with $9$ ordinary double points and no other singularities.
Proposition 4.16.
$S_{3}$ admits a fibration with no section, singular fibre types $I_{2}^{*}{I_{0}^{*}}^{2}I_{2}^{2}$, full level-$2$ structure, and fibres of degree $6$. The Jacobian admits a $2$-isogeny to the elliptic surface $\mathcal{E}^{\prime}$ defined by
$y^{2}=x^{3}+(-81t^{3}-95t^{2}+176t)x^{2}+(256t^{4}-512t^{3}+256t^{2})x,$
which has a point with $x$-coordinate $(4t+9t^{2}/2)^{2}$ and height $3$.
Proof.
Again this is done computationally. The only slightly tricky point is that, though the generic fibre has degree $6$ and arithmetic genus $2$, the projection away from the singular point is a map to a rational curve, not to a singular quartic. Thus, to determine the Weierstrass model, we need to start by using the fibre and its Veronese embedding to place the fibre in $\mathbb{P}^{9}$. This done, we can project away from singular points until we reach an intersection of two quadrics in $\mathbb{P}^{3}$. ∎
Remark 4.17.
The fibration on $S_{3}$ was chosen precisely because it is $2$-isogenous to a fibration with $I_{4}^{*}{I_{0}^{*}}^{2}I_{1}^{2}$ fibres, which matches one of the types of fibration on the Kummer surface of $E\times E^{\prime}$. Thus we know that, at least over $\bar{\mathbb{Q}}$, the surface $\mathcal{E}^{\prime}$ is isomorphic the Kummer surface of a product of two elliptic curves.
In order to facilitate the comparison between $\mathcal{E}$ and the Kummer surface, we now find the fibration with reducible fibre types $I_{6}^{*}I_{2}^{6}$. This is again done as in Section 2.2: we find a model in $\mathbb{P}^{5}$ with at worst $A_{3}$ singularities and find such a fibration there. We conclude:
Proposition 4.18.
There is a fibration on $S_{3}$ whose generic fibre has reducible fibres $I_{6}^{*},I_{2}^{6}$ which is defined by the equation
$y^{2}=x^{3}+(6t^{3}-119/4t^{2}-128t+147/4)x^{2}+(-54t^{3}+1071/4t^{2}+1152t-1647/4)x.$
We now start again from the other side, with the Kummer surface of $E\times E^{\prime}$ where $E,E^{\prime}$ are generic elliptic curves with full level-$2$ structure defined by equations $y^{2}=x(x-1)(x-p)$ and $y^{2}=x(x-1)(x-p^{\prime})$ where $p,p^{\prime}$ are independent transcendentals. (These calculations use code originally written for [20] and my coauthor Jared Weinstein had considerable influence on it.) We find the equation of a fibration with $I_{6}^{*}I_{2}^{6}$ fibres and identify $p$ and $q$ by matching the locations of the $I_{2}$ fibres. In particular, we choose $p,p^{\prime}=-87\pm 24\sqrt{13}$ to find two elliptic curves $E,E^{\prime}$ of conductor $(3)$ over ${\mathbb{Q}}(\sqrt{13})$ that are $3$-isogenous. It is easy to write down a model over ${\mathbb{Q}}$ for the Kummer surface of the product and the $I_{6}^{*}I_{2}^{6}$ fibration on this surface. However, it cannot be isomorphic to that on the other one, because the dimensions of the Galois-fixed subspaces are different: $11$ on the Kummer surface and $16$ on the image of the Hilbert modular surface. As before, we need to take the Jacobian of a fibration with no rational sections. We state the result as a proposition.
Proposition 4.19.
There is a fibration on the Kummer surface of $E\times E^{\prime}$ defined over ${\mathbb{Q}}$ that has a $\bar{\mathbb{Q}}$-section but no rational section. Its Jacobian is defined by the Weierstrass equation
$y^{2}=x^{3}+(6t^{3}-(119/4)t^{2}-128t+(147/4))x^{2}+(-54t^{3}+(1071/4)t^{2}+1152t-(1647/4))x.$
Remark 4.20.
It was sensible to look for this particular fibration, because taking its Jacobian over ${\mathbb{Q}}$ increased the rank of the rational part of the Picard group substantially. An $I_{6}^{*}$ fibre may only contribute rank $5$ to the rational part of the Picard lattice when there is no rational section, but must contribute at least $10$ when there is one, because the rational section forces the component it meets to be rational, and therefore all components except for the two farthest from the zero component to be.
There is an ambiguity when we speak of $\mathcal{K}$ as a K3 surface over ${\mathbb{Q}}$, since we could twist by $13$ to obtain another K3 surface whose base change to ${\mathbb{Q}}(\sqrt{13})$ is isomorphic to that of $\mathcal{K}$. It is easy to see which one must be correct, since the number of points mod $p$ must match mod $p$ for all primes $p$ of good reduction.
We observe that the Weierstrass equations in Propositions 4.18 and 4.19 are identical. The following theorem summarizes the content of this section:
Theorem 4.21.
Let $H_{I,K}$ be the Hilbert modular surface of level $(3)$ over ${\mathbb{Q}}(\sqrt{13})$ and let $\mathcal{K}$ be the Kummer surface of $E\times E^{\sigma}$, where $E$ is the elliptic curve defined by $y^{2}=x(x-1)(x-(87+24\sqrt{13}))$. Then there is a K3 surface $S$ with maps of degree $32,4$ from $H_{I,K}$ and $\mathcal{K}$ respectively, such that the correspondence associated to these maps induces an isogeny of the transcendental lattices of $H_{I,K}$ and $\mathcal{K}$.
In other words, we have made the Oda-Hamahata conjecture completely explicit in this case.
4.4. Level ${\mathfrak{p}}_{2}^{3}$ over ${\mathbb{Q}}(\sqrt{17})$
We now consider the case of level ${\mathfrak{p}}_{2}^{3}$ over ${\mathbb{Q}}(\sqrt{17})$, where ${\mathfrak{p}}_{2}$ is a prime of norm $2$. In this case the level is not Galois-stable and so we do not have a ${\mathbb{Q}}$-curve; the conjecture was not already known in this case.
In this example we choose not to use the Atkin-Lehner quotient. The ring of Hilbert modular forms is generated by $5$ forms of weight $2$ and $7$ of weight $4$ and thus its Proj embeds into $\mathbb{P}(1^{5},2^{7})$ and hence into $\mathbb{P}^{21}$ by ${\mathcal{O}}(2)$. We find that the image $S_{21}$ is defined by $151$ quadrics and is of degree $64$.
Proposition 4.22.
There is a birational map from $S_{21}$ to a surface $S_{5}^{\prime}$ in $\mathbb{P}^{5}$ defined by $3$ quadrics and whose only singularities are $6$ ordinary double points.
Proof.
We can find $4$ singular points on $S_{21}$, as we should since there are $4$ cusps for $\Gamma_{0}(8)$ over ${\mathbb{Z}}$ and no elliptic points; the projection away from these gives a map to $\mathbb{P}^{17}$ whose image $S_{17}$ is a surface of degree $44$ which again has $4$ singular points. Projecting away from these we now have a surface $S_{13}$ of degree $36$ in $\mathbb{P}^{13}$ and yet again $4$ singular points. This time the projection goes to $S_{9}\subset\mathbb{P}^{9}$, which has degree $24$ and (at least) $7$ singular points, and by a careful choice of which four to project from we obtain a surface $S_{5}$ of degree $12$ in $\mathbb{P}^{5}$. A final projection from $S_{5}$ yields a surface $S_{4}$ of degree $10$ in $\mathbb{P}^{4}$.
This surface in $\mathbb{P}^{4}$, in turn, has singularities along two lines and three points (pairwise disjoint), and the linear system of quadrics vanishing on these gives a map to a different singular model $S_{4}^{\prime}\subset\mathbb{P}^{4}$, this time of degree $8$. The linear system of quadrics vanishing on the two singular lines and two singular points of this surface gives a map to a surface $S_{5}^{\prime}\subset\mathbb{P}^{5}$ as described in the statement. ∎
The surface $S_{5}^{\prime}$ is a K3 surface. One way to find genus-$1$ fibrations on a K3 surface in $\mathbb{P}^{n}$ is to consider the families of hyperplanes through $n-1$ singular points. In this case it turns out that one of the $\binom{6}{4}$ sets of singular points is coplanar, but the other $14$ do indeed give genus-$1$ fibrations. In no case does the degree of the fibre exceed $4$, so Magma’s built-in functions easily find the Jacobians of these fibrations and hence their reducible fibres. It turns out that all $44$ components of these are defined over ${\mathbb{Q}}(\sqrt{17})$ and that, together with the curves above the nodes, they generate a sublattice of the Picard lattice of rank $18$ and discriminant $-64$. We can verify that the rank is $18$ by reducing modulo small primes and finding an irreducible factor of degree $4$ in the $L$-function. We also verify that we have found the full Picard lattice as follows.
Proposition 4.23.
Let $S$ be a K3 surface with an elliptic fibration $\pi$ with reducible fibres $A_{7},A_{1}^{8}$ and rank $1$. Then the discriminant of $\operatorname{Pic}S$ cannot be $1$, $4$, or $16$.
Proof.
According to the tables of [30], these reducible fibres force the fibration to have full $2$-torsion, which means that if the generator $s$ of the Mordell-Weil group modulo torsion has height $h$ then the discriminant is $(8\cdot 2^{8}\cdot h)/2^{4}$. Thus we would need $h\in\{1/128,1/32,1/8\}$. By the results of [9], the first two are impossible: the height is the sum of an integer with numbers of the form $a(8-a)/8,b(2-b)/2$, so there is no way for the denominator to be greater than $8$. There are several ways to get height $1/8$, though, and we must exclude them by showing that they are inconsistent with $2$-torsion: in particular, by proving that the height of such a section would not be the same as that of its $2$-torsion translates. Note that there are two $2$-torsion sections that pass through component $4$ of the $A_{7}$ fibre and the nonzero component of four of the $A_{1}$’s and one that meets the zero component of the $A_{7}$ and the nonzero components of all $A_{1}$’s.
A more careful examination of [9] shows that there are three possibilities for $s$:
  1. (1)
    The section $s$ meets the zero section once, the nonzero components of all $A_{1}$ fibres, and a component of $A_{7}$ three away from the origin (so that its height is $4+2-8\cdot(1/2)-3\cdot 5/8=1/8$). Translating this by the $2$-torsion section that meets the nonzero components of all $A_{1}$ gives a section of height $84+2e-3\cdot 5/$ 8, where $e$ is the intersection with the zero section. Since $e$ cannot be $-1$ this is not $1/8$.
  2. (2)
    The section $s$ meets a component of $A_{7}$ three away from the origin and four nonzero components of $A_{1}$ fibres. If we translate by a torsion point passing through the far component of the $A_{7}$, the height becomes $24+2e-7/8-n/$ 2, where $n$ is the number of nonzero components of $A_{1}$ fibres now met, so we must have $e=0,n=6$. However, this holds for both such torsion points, which is impossible because their difference passes through the nonzero component of every $A_{1}$.
  3. (3)
    The section $s$ meets a component of $A_{7}$ one away from the zero component and the nonzero components of six $A_{1}$ fibres. This is excluded by the same argument as in the first case.
Corollary 4.24.
The discriminant of the Picard lattice of $S_{5}^{\prime}$ is $-64$.
Proof.
Since we have found a sublattice of maximal rank and discriminant $-64$, the discriminant of the full lattice is $-64/n^{2}$ for some integer $n$. One of the $14$ fibrations defined from $3$-planes through $4$ singular points as described above does indeed have reducible fibres as in Proposition 4.23, which shows that $n^{2}=1$. ∎
Listing $ADE$ configurations among the $50$ known curves, we look for those that have multisection degree $2$ and have fibre class $F$ such that enlarging the lattice by $F/2$ produces a lattice isomorphic to that of the Kummer surface of $E\times E^{\prime}$ (which can be described as $D_{8}\oplus D_{8}\oplus U$, where $U$ is the hyperbolic lattice $\langle x,y\rangle$ with $x^{2}=y^{2}=0,x\cdot y=1$, or in various other ways).
Proposition 4.25.
There is a fibration on $S_{5}^{\prime}$ with multisection degree $2$, two reducible $I_{8}$ fibres, and fibres of degree $8$.
Proof.
This is easily calculated from knowledge of the Picard lattice and the curves we have found on $S_{5}^{\prime}$. One way to find its equations is to write down the two reducible fibres $F_{1},F_{2}$ as subschemes of $S_{5}^{\prime}$. Let $Q_{1}$ be a quadratic polynomial vanishing on $F_{1}$ but not all of $S_{5}^{\prime}$, so that $S_{5}^{\prime}\cap(Q_{1}=0)=F_{1}\cup R$. Then $F_{2}\cup R$ is also cut out by a quadratic polynomial $Q_{2}$, and the map given by $(Q_{1}:Q_{2})$ is the desired genus-$1$ fibration, though with base locus $R$. ∎
Projecting away from a suitable singularity, we express the generic fibre of this fibration as a curve of degree $6$ in $\mathbb{P}^{4}$, and from there by projecting away from a divisor of degree $2$ we get a singular quartic. In contrast to the slightly more complicated situation of Proposition 4.16, this allows us to find a smooth model in $\mathbb{P}^{3}$ by mapping by the linear system of quadrics through the singular points. We can then find the Jacobian of this fibration (call it $J$).
Proposition 4.26.
$J$ is isomorphic to the elliptic curve defined by $y^{2}=x^{3}+(t^{4}+13t^{2}+16)x^{2}+4t^{4}x$.
The curve thus defined spreads out to a K3 surface with the Picard lattice of a Kummer surface of the product of two elliptic curves. Using the calculations of [20], we compare this fibration to the fibrations on the Jacobian of a product of two elliptic curves with full level-$2$ structure, and we indeed find a match with $p=(-5a-4)/4$, where $a^{2}-a-4=0$ is a generator of the field ${\mathbb{Q}}(\sqrt{17})$. This produces one of the curves of conductor ${\mathfrak{p}}_{2}^{3}$, where ${\mathfrak{p}}_{2}$ is one of the primes above $2$ in ${\mathbb{Q}}(a)$. Let $q=p^{\sigma}$ where $\sigma$ is the nontrivial automorphism of ${\mathbb{Q}}(a)$.
Starting again from the other side, we take the elliptic curves $E:y^{2}=x(x-1)(x-p),E^{\sigma}:y^{2}=x(x-1)(x-q)$ and consider the Kummer surface of their product. To construct these curves from the given ones whose conductor has norm $8$, we need to take a quadratic twist by an element whose norm is the negative of a square, so the Kummer surface we are actually interested in is a quadratic twist of this one by $-1$. (We use $E\times E^{\sigma}$ because the formulas for its fibrations are already available in a convenient form.) Since $p$ and $q$ are quadratic conjugates, it is easy to find a model of this surface that is defined over ${\mathbb{Q}}$ and to pull back the fibrations with two $I_{8}$ fibres to it. Only one of them has bad fibres in locations that match those of $J$. It has fibres that are defined over ${\mathbb{Q}}$, so it is a simple matter to rewrite this fibration in terms of equations defined over ${\mathbb{Q}}$. As in Section 4.3, the surfaces are isomorphic over ${\mathbb{Q}}$ up to a possible twist by $17$ and we need to choose the correct ${\mathbb{Q}}$-model of the Kummer surface to obtain an isomorphism over ${\mathbb{Q}}$.
4.5. Level $(2)$ over ${\mathbb{Q}}(\sqrt{17})$
In this section we consider a case where the level is Galois stable and thus the conjecture is known. Nevertheless there are some interesting questions to study.
Since $2$ splits in ${\mathbb{Q}}(\sqrt{17})$, there are $4$ cusps and hence $4$ Eisenstein series in weight $2$. It turns out that the ring of modular forms is generated by these together with a cusp form of weight $2$ and $3$ cusp forms of weight $4$.
Proposition 4.27.
The quotient of the Hilbert modular surface $H_{(2),{\mathbb{Q}}(\sqrt{17})}$ by the Atkin-Lehner involutions is birational to a K3 surface of degree $6$ in $\mathbb{P}^{4}$ with singularities of type $A_{1},A_{2},A_{3}$.
Proof.
Since the cusp form is fixed by both Atkin-Lehner involutions, we immediately pass to the quotient by these, obtaining a surface in $\mathbb{P}^{5}$. This surface is singular along a line, away from which we project to create a surface in $\mathbb{P}^{3}$ singular along a conic. The quadratic forms vanishing on this conic give a projectively normal embedding of a K3 surface into $\mathbb{P}^{4}$ with the usual degree $6$. The singularities are easily checked in Magma. ∎
Projecting away from the unique $A_{3}$ gives a new model $S_{3}\subset\mathbb{P}^{3}$ with a rational $A_{2}$ singularity, two rational $A_{1}$s, and four pairs of $A_{1}$ singularities defined and conjugate over ${\mathbb{Q}}(\sqrt{2})$. The outer components of the $A_{3}$ are now lines over ${\mathbb{Q}}(\sqrt{17})$. Using this information we are able to determine the entire Picard group. It has rank $19$; this is as expected, because the elliptic curves over ${\mathbb{Q}}(\sqrt{17})$ of conductor $(2)$ are ${\mathbb{Q}}$-curves. The determinant of the Gram matrix is $192$ and the generators of the Picard group are all defined over ${\mathbb{Q}}(\sqrt{2},\sqrt{17})$.
Since some of the curves of conductor $(2)$ are $2$-isogenous to their Galois conjugates, there should be a correspondence with the Kummer surface of a product $E\times E^{\prime}$, where $E,E^{\prime}$ are $2$-isogenous. Such a surface has Picard discriminant $32$. Other curves in the isogeny class are $8$-isogenous to their conjugates, resulting in a more complicated Kummer surface. Our considerations in this paper are only up to maps of finite degree, so we need not treat these.
Let us first describe how we found a correspondence between the Hilbert modular surface and the Kummer surface of $E\times E^{\prime}$ defined over ${\mathbb{Q}}(\sqrt{17})$. As above, we construct a model $S_{3}$ of the Atkin-Lehner quotient of $H_{(2),{\mathbb{Q}}(\sqrt{17})}$ as a K3 surface in $\mathbb{P}^{3}$. By standard calculations we find:
Proposition 4.28.
There is a genus-$1$ fibration on $S_{3}$ with $\bar{\mathbb{Q}}$-multisection degree $2$ and ${\mathbb{Q}}$-multisection degree $4$. The reducible fibre types are $I_{2}^{*},I_{1}^{*},I_{4},I_{2},I_{2}$, and a Weierstrass equation for the Jacobian is
$y^{2}$ $=x^{3}+(-2t^{3}-359/16t^{2}-99/2t+33)x^{2}+$
$\qquad(t^{6}+55/2t^{5}+871/4t^{4}+4937/8t^{3}+523/2t^{2}-842t+264)x.$
The Jacobian admits a model $S_{7}\subset\mathbb{P}^{7}$ with $8$ $A_{1}$ singularities, all defined over ${\mathbb{Q}}$, and no others; its Picard group has discriminant $48$ and the generators are all defined over ${\mathbb{Q}}(\sqrt{17})$. It has a fibration with bad fibre types $I_{2}^{*},I_{2}^{*},I_{0}^{*}$ and a $2$-torsion section.
The importance of this fibration is that vertical curves together with the sections of order $1$ and $2$ contain a configuration that is a fibre of a fibration with reducible fibres $I_{2}^{*},{I_{0}^{*}}^{2},I_{2}^{2}$ and full level-$2$ structure. One of the $2$-torsion sections must pass through the reduced component of the $D_{6}$ that is on the same side as the zero section; the formulas of [9] imply that, in order to have height 00, it must pass through the nonzero components of the $A_{1}$ fibres. Accordingly the quotient by this $2$-isogeny will have reducible fibres $D_{8},D_{4}^{2}$. Every K3 surface with a $D_{8},D_{4}^{2}$ fibration is the Kummer surface of a product of elliptic curves.
Unfortunately not all of the components of the fibres of the fibration on $S_{7}$ are rational. Although we believe that there should be a correspondence over ${\mathbb{Q}}$, we have not been able to exhibit it, for reasons to be discussed in Remark 4.30. We therefore work over ${\mathbb{Q}}(\sqrt{17})$ for the rest of this section.
Having constructed the ${I_{2}^{*}}^{2},I_{0}^{*}$ fibration as an elliptic curve over ${\mathbb{Q}}(\sqrt{17})(u)$, we use a model in $\mathbb{P}^{7}$ to help construct the desired $I_{2},{I_{0}^{*}}^{2},I_{2}^{2}$ fibration. Magma immediately produces the $2$-isogenous elliptic surface with $I_{4}^{*},{I_{0}^{*}}^{2}$ fibres and a point of order $2$.
Since the Picard discriminant was $48$ before the last $2$-isogeny, the Mordell-Weil generator must have been of height $3$. It cannot be of height $3/2$ on the new surface, since by [9] every section on a surface whose reducible fibres are all of type $I_{2n}^{*}$ must have integral height; thus the height must be $6$ and we expect that the surface thus obtained will be the Kummer surface of a product of $6$-isogenous elliptic curves.
Theorem 4.29.
The elliptic surface we have constructed with $I_{4}^{*},{I_{0}^{*}}^{2}$ fibres is isomorphic to the Kummer surface of 2.2.17.1/4.1-a/8 $\times$ 2.2.17.1/4.1-a/10.
Proof.
We write down a suitable fibration on the product using the formulas from [20]. That is, by writing down the $D_{8},D_{4}^{2}$ fibrations on the Kummer surface of a general $E\times E^{\prime}$ with full level-$2$ structure we see how to choose the parameters in the Legendre family so that the local invariants match and then we verify that the fibrations are exactly isomorphic. ∎
Strictly speaking, the Oda-Hamahata conjecture predicts a correspondence with the Kummer surface of $E\times E^{\sigma}$, but an isogeny from $E^{\prime}$ to $E^{\sigma}$ induces one from $\mathcal{K}(E\times E^{\prime})$ to $\mathcal{K}(E\times E^{\sigma})$.
Remark 4.30.
In attempting to construct a correspondence over ${\mathbb{Q}}$, the basic problem is that the curves in isogeny class 2.2.17.1/4.1-a are $2$- or $8$-isogenous to their Galois conjugates and so the Picard discriminants of the obvious Kummer surfaces to consider over ${\mathbb{Q}}$ are not multiples of $3$. Starting from the surface of Picard discriminant $192$, therefore, we must at some point find a $3$-isogeny to make the $3$-adic valuation of the discriminant even. While this is not difficult to do, it always resulted in a surface whose Picard discriminant was divisible by $9$. To pass to the desired Kummer surface would then require taking a $3$-Jacobian of a fibration with no section. Again, it is easy to find such a fibration, but although many surfaces were investigated in the hope that one of them would have such a fibration defined over ${\mathbb{Q}}$, none had this property. Thus the desired correspondence could not be constructed over ${\mathbb{Q}}$.
4.6. Level $(3)$ over ${\mathbb{Q}}(\sqrt{2})$
This example is somewhat similar to the previous one. The elliptic curves are in the isogeny class 2.2.8.1/9.1-a; they are curves defined over ${\mathbb{Q}}$ twisted by $2+\sqrt{2}$. Since $N_{{\mathbb{Q}}(\sqrt{2})/{\mathbb{Q}}}(2+\sqrt{2})$ is a square in ${\mathbb{Q}}(\sqrt{2})$, this means that they are isomorphic to their Galois conjugates. The curves do not have complex multiplication, so the Kummer surface of $E\times E^{\sigma}$ will have geometric Picard number $19$.
The ring of modular forms is generated by forms of weight $2,2,2,4,6,6,8$; there are two Eisenstein series of weight $2$, and all the rest of the generators may be taken to be cusp forms.
Lemma 4.31.
There is a birational map from the Atkin-Lehner quotient to a K3 surface in $\mathbb{P}^{9}$. This surface admits a genus-$1$ fibration whose Jacobian has Weierstrass equation
$y^{2}=x^{3}+(2t^{4}+3t^{3}-7t^{2}+6t+14)x^{2}+(18t^{5}+45t^{4}-36t^{3}-90t^{2}+18t+45)x$
with reducible fibre types $I_{0}^{*},I_{6},I_{4},I_{2}^{4}$, full level-$2$ structure, and a section of height $5/12$ defined over ${\mathbb{Q}}(\sqrt{2})$ whose $x$-coordinate is $66t^{2}-$ 6. The discriminant of the Picard lattice is $40$.
Proof.
Since $3$ is inert in ${\mathbb{Q}}(\sqrt{2})$, there is a unique Atkin-Lehner operator. We consider the map given by the sections of ${\mathcal{O}}(4)$ in its $+1$ eigenspace; these give a map to $\mathbb{P}^{12}$ whose image we compute by Algorithm 2.7 to be a surface of degree $24$. The projection away from the image of the cusps is of degree $22$ and is singular along a line; projecting away from this line produces a surface of degree $16$ in $\mathbb{P}^{9}$ defined by quadrics. Using Magma we check that this surface has trivial canonical sheaf and then that it is a K3 surface. On this surface we are fortunate to be able to find some curves of genus $1$ by setting one of the coordinates to be 00. That is, some of the components of the subscheme of $S_{9}$ defined by $x_{i}=0$ turn out to be curves of arithmetic genus $1$ and degree $4$. We find the fibrations of which these curves are fibres and pass to the Jacobian; for one of them, the reducible fibre types are as described and as in Remark 3.5 we find the given Weierstrass equation. The section may be found by standard techniques, although in this example it is easy enough to spot it by reducing mod small $p$ and searching.
The calculation of the discriminant follows from the determinant formula [29, Corollary 6.39] as before: it is $404\cdot 6\cdot 4\cdot 2^{4}\cdot 5/12\cdot(1/4)^{2}=$ 40. ∎
Remark 4.32.
Note that, as in the previous example, there is a prime $p>2$ dividing the Picard discriminant which is the isogeny degree between some curves in this isogeny class but does not divide the isogeny degree between a curve and its Galois conjugate. Thus, in order to find the desired correspondence over ${\mathbb{Q}}$, we would at some point need to use a map of degree divisible by $5$. In fact the Picard lattice is isomorphic to that of a K3 surface with $4$ fibres of type $I_{5}$, a section of height $8/5=4-2(6/5)$, and a $5$-torsion section. The $5$-isogeny goes to a surface of Picard discriminant $200$; however, as before it seems impossible to find a genus-$1$ fibration on this surface defined over ${\mathbb{Q}}$ with multisection degree $5$, so we cannot find the correspondence with a Kummer surface over ${\mathbb{Q}}$.
We proceed to find a correspondence with a Kummer surface over ${\mathbb{Q}}(\sqrt{2},\sqrt{3})$, the $2$-division field of two of the curves in the isogeny class.
Proposition 4.33.
The elliptic surface of Lemma 4.31 admits a model in $\mathbb{P}^{3}$ with three rational $A_{2}$ singularities and one pair each of conjugate $A_{2}$ and $A_{1}$. It has a fibration with reducible fibre types $I_{2}^{*},{I_{0}^{*}}^{2},I_{2}^{2}$ and full level-$2$ structure. One of the curves $2$-isogenous to this one has reducible fibres $I_{4}^{*},{I_{0}^{*}}^{2}$ and a point of order $2$. The equation of the generic fibre of this fibration is
$y^{2}=x^{3}+(-10t^{3}-88t^{2}+98t)x^{2}+(t^{6}-4t^{5}+6t^{4}-4t^{3}+t^{2})x.$
Proof.
As before, this is a standard calculation for which the reader is referred to [21]. ∎
Remark 4.34.
For the reason to seek out this type of fibration, compare Section 4.3; the point is that the isogenous fibration is found on the Kummer surface of a product.
As before, the formulas from [20] identify this fibration over ${\mathbb{Q}}(\sqrt{2},\sqrt{3})$ with one on the Kummer surface of $y_{0}^{2}-x_{0}(x_{0}-1)(x_{0}-p)=y_{1}^{2}-x_{1}(x_{1}-1)(x_{1}-q)$ where $p=6+2\sqrt{6},q=-485+198\sqrt{6}$. These are twists of the curves 2.2.8.1/9.1-a/3, 2.2.8.1/9.1-a/4 by the same number $-2+\sqrt{2}-\sqrt{3}+\sqrt{6}$, so the Kummer surfaces of the two products are isomorphic over ${\mathbb{Q}}(\sqrt{2},\sqrt{3})$. We thus conclude:
Corollary 4.35.
Over ${\mathbb{Q}}(\sqrt{2},\sqrt{3})$ there is a correspondence between the Hilbert modular surface $H_{(3),{\mathbb{Q}}(\sqrt{2})}$ and the product of an elliptic curve over ${\mathbb{Q}}(\sqrt{2})$ of conductor $(3)$ and its conjugate.
Proof.
Above we constructed a dominant map from $H_{(3),{\mathbb{Q}}(\sqrt{2})}$ to one such Kummer surface. As before, any two such Kummer surfaces admit maps induced by isogenies of the corresponding abelian varieties. ∎
4.7. Level ${\mathfrak{p}}_{17}$ over ${\mathbb{Q}}(\sqrt{2})$
In this case, where again the conjecture was not previously known, we are able to give a correspondence over ${\mathbb{Q}}(\zeta_{8})$. In this section we let $K={\mathbb{Q}}(\sqrt{2})$ and $I=(5-2\sqrt{2})$, a prime ideal of ${\mathcal{O}}_{K}$ of norm $17$. The Hilbert modular variety $H_{I,K}$ is a subvariety of $\mathbb{P}(1^{3},2^{4})$.
Proposition 4.36.
The image of the Atkin-Lehner quotient of $H_{I,K}$ by ${\mathcal{O}}(2)$ is a surface in $\mathbb{P}^{5}$ of degree $12$ defined by equations of degree $2,3,3,4,4,4,4,4,4$ and is birational to a surface with a genus-$1$ fibration whose Jacobian is a K3 surface. It has two conjugate $I_{0}^{*}$ fibres defined over ${\mathbb{Q}}(\sqrt{17})$ as well as one $I_{4}$ and three $I_{2}$ fibres defined over ${\mathbb{Q}}$. There are two sections defined over ${\mathbb{Q}}(\sqrt{2})$ of height $3/4$ and $3/2$ and height pairing 00. The discriminant of the Picard lattice is $-144$.
Proof.
The first statement is a routine calculation. The image of the cusps is a singularity $C$ such that the projection away from $C$ has degree $8$ in $\mathbb{P}^{4}$. Considering forms of degree $2$ vanishing along the line where this surface is singular and the worst of the singular points, we map to a surface $S_{9}$ of degree $21$ in $\mathbb{P}^{9}$. Now $S_{9}$ is singular along two lines; projecting away from their span, we find a surface of degree $10$ in $S_{5}$ defined by polynomials of degree $2,2,3,3$. Although this is still singular on a line, we can find a genus-$1$ fibration on it (defined by linear forms and with base scheme consisting of the singular line doubled and a certain singular curve of degree $4$). The Jacobian of this fibration is an elliptic surface with Euler characteristic $24$, which is therefore a K3 surface. Once the equation is found, the reducible fibres are easy to check and the sections can be found be standard techniques.
Given this information, we verify that the subgroup of the Mordell-Weil group that we have found is saturated at $2$ and $3$. Since the torsion subgroup has order $2$, it follows that the Picard discriminant is $-(4^{2}\cdot 4\cdot 2^{3}\cdot 3/2\cdot 3/4)/2^{2}=-144$. ∎
Remark 4.37.
The minimal desingularization of $S_{5}$ has the same Hodge diamond as a K3 surface, but it is not itself a K3 surface. This follows, for example, from Kodaira’s formula for the canonical bundle of a surface with a genus-$1$ fibration, combined with the existence of a point of $\mathbb{P}^{1}$ where the fibre is nonreduced on $S_{5}$ but is a smooth rational curve on the Jacobian. The situation is analogous to that of an Enriques surface, for which the Hodge diamond is the same as that of a rational elliptic surface.
A somewhat similar calculation exhibits a genus $1$ fibration on $H_{I,K}$; alternatively one can verify that the ramification locus of the quotient map is disjoint from the generic fibre of the fibration. As mentioned in Remark 4.1, this shows that Hamahata’s claim that $\kappa(H_{I,K})=2$ is incorrect.
Having determined a K3 surface quotient of the Hilbert modular surface, we must now improve the model. As usual, we do this by embedding in $\mathbb{P}^{5}$ and successively projecting away from noncanonical singularities and applying the Veronese embedding until the desired model is reached. In this case the procedure is quite lengthy, requiring $10$ projections, but once this is done we can project away from the unique $A_{3}$ point to find a model $S_{7}\subset\mathbb{P}^{7}$ whose only singularities are $11$ ordinary double points. The fibres of the fibration are of degree $3$. In this model, the reducible fibres are two of type $I_{0}^{*}$, defined over ${\mathbb{Q}}(\sqrt{17})$ and given in this model by a double line, a line, and three rational nodes; one $I_{4}$, consisting of a rational line, two conjugate lines over ${\mathbb{Q}}(\sqrt{2})$, and a rational node; and three $I_{2}$, whose components are a node and a cubic singular there.
We might hope to reach a Kummer surface of a product by taking the Jacobian of a fibration with a $3$-section but no rational section. There are two problems with this: first, no rational Picard class that is not a multiple of $3$ has intersection divisible by $3$ with all rational Picard classes, and second, the lattice containing the Picard lattice with index $3$ is not the Picard lattice of the Kummer surface of a product of two elliptic curves but rather another lattice of the same rank and discriminant.
Proposition 4.38.
There is a fibration on $S_{7}$ with a $2$-section but no section. It has reducible fibres $I_{4}^{*},I_{6},I_{2},I_{2}$, torsion of order $2$, and a section of height $3/2$. The Jacobian has Weierstrass model
$y^{2}=x^{3}+(17t^{3}-14t^{2}+7t-2)x^{2}+(4t^{4}-13t^{3}+15t^{2}-7t+1)x,$
and the $x$-coordinate of a section is $1$.
We proceed as before to find a good projective model of this surface. In this case it is easy to find a K3 surface model, but its singularities are undesirable and we have to do some work to find a model $S_{6}\subset\mathbb{P}^{6}$ with singularities of type $A_{3},A_{3},A_{2},A_{1}$. It turns out that the full Picard lattice is defined over ${\mathbb{Q}}(\sqrt{2})$; we no longer need to adjoin $\sqrt{17}$.
Unfortunately there is still no rational divisor class that can be divided by $3$; this remains true for other isogenous surfaces, so we consider a genus-$1$ fibration defined over ${\mathbb{Q}}(\sqrt{2})$ with multisection degree $3$.
Proposition 4.39.
There is a fibration on $S_{6}$ defined over ${\mathbb{Q}}(\sqrt{2})$ with two $I_{4}^{*}$ fibres, torsion of order $2$, and multisection $3$.
Again, the reason for wanting such a fibration is that the $2$-isogenous surface has two $I_{2}^{*}$ fibres, four of type $I_{2}$, and $2$-torsion, and this is one of the type of fibrations on the Kummer surface of $E\times E^{\prime}$.
The rest is now routine; as in the previous examples, we use the results of [20] to match this fibration with one on the Kummer surface of $E\times E^{\prime}$. It is necessary to extend the field from ${\mathbb{Q}}(\sqrt{2})$ to ${\mathbb{Q}}(\zeta_{8})$ where the $2$-torsion of $E,E^{\prime}$ is defined, but in the end we prove the following result:
Theorem 4.40.
The Hilbert modular surface $H_{I,K}$ admits a rational map of finite degree defined over ${\mathbb{Q}}(\zeta_{8})$ to the Kummer surface of $E\times E^{\prime}$, where $E,E^{\prime}$ are elliptic curves over $K$ of conductor $I$.
As before, if this is true for one choice of $E,E^{\prime}$ it is true for all of them.
4.8. Level ${\mathfrak{p}}_{31}$ over ${\mathbb{Q}}(\sqrt{5})$
The Kodaira dimension of the Hilbert modular surface corresponding to the ideal of smallest norm in ${\mathcal{O}}_{{\mathbb{Q}}(\sqrt{5})}$ giving a surface with $p_{g}=1$ was left undecided by Hamahata. Here we will prove that its Kodaira dimension is $1$ and verify the Oda-Hamahata conjecture. We begin as usual by computing the ring of Hilbert modular forms, finding it to be generated by $3,6,4$ forms of weight $2,6,10$ respectively. We also determine the Atkin-Lehner involution and verify that it has $8$ isolated fixed points, none of which is singular on the surface.
Proposition 4.41.
The Atkin-Lehner quotient of the Hilbert modular surface $H_{(6-\sqrt{5}),{\mathbb{Q}}(\sqrt{5})}$ admits a fibration by curves of genus $1$ whose Jacobian is a K3 surface.
Proof.
We begin by studying the quotient. There are too many modular forms of weight $10$, so we use the forms of weight $6$ invariant under Atkin-Lehner to map to a surface of degree $18$ in $\mathbb{P}^{8}$. By projecting successively from the cusp, the new singular point in its image, and the singular line we reach a surface of degree $9$ in $\mathbb{P}^{4}$. Although this surface does not have canonical singularities, it is sufficiently well-behaved that Magma can compute a canonical divisor map whose codomain is $\mathbb{P}^{1}$ and whose general fibre is of genus $1$, verifying that the surface is not of general type. ∎
The surface has one $I_{3}$, ten $I_{2}$, and one $I_{1}$ fibre, indicating that its Jacobian is a K3 surface. However, we claim that its Kodaira dimension is $1$.
Proposition 4.42.
The quotient of $H_{{\mathfrak{p}}_{31},{\mathbb{Q}}(\sqrt{5})}$ by the Atkin-Lehner involution has Kodaira dimension $1$.
Proof.
Since the Euler characteristic is $24$, this is equivalent to the statement that the fibration has at least one multiple fibre. This is easy to verify computationally ∎
Corollary 4.43.
The Hilbert modular surface $H_{{\mathfrak{p}}_{31},{\mathbb{Q}}(\sqrt{5})}$ has Kodaira dimension $1$.
Proof.
Let $F$ be the fixed locus of the Atkin-Lehner involution. Since $F$ consists of isolated nonsingular points, its image on the Atkin-Lehner quotient consists of ordinary double points. Thus, on the minimal resolution of the quotient, the double cover is ramified precisely at the $8$ exceptional curves of self-intersection $-2$ above these points. Since the fibration on the quotient is well-defined at these points, the curves are vertical for the fibration, and hence the general fibre of the genus-$1$ fibration on the quotient is disjoint from the ramification locus. It follows that the general fibre pulls back to a curve of genus $1$. Thus $H_{{\mathfrak{p}}_{31},{\mathbb{Q}}(\sqrt{5})}$ admits a genus-$1$ fibration and cannot be of general type. ∎
We now proceed to exhibit a correspondence between the Jacobian of the Atkin-Lehner quotient and the Kummer surface of the product of an elliptic curve in the isogeny class and its conjugate. This is a relatively easy example because the diameter (Definition 4.5) of the isogeny class is $4$ (cf. Claim 4.6) and there are curves in the class with full level-$2$ structure. As usual, we do most of the work on the side of the quotient rather than that of the Kummer surface.
The eight ordinary double points mentioned above are components of $I_{2}$ fibres and are defined over a number field whose Galois group has order $64$. It would be unpleasant to work with a K3 surface over such a large number field, so we begin by passing to the $2$-isogenous elliptic surface $\mathcal{E}$ on which the images of these fibres are of type $I_{1}$. The reducible fibres of $\mathcal{E}$ are of type $I_{6},I_{4},I_{4},I_{2}$.
Proposition 4.44.
The Weierstrass model of $E$ can be taken to be
$y^{2}$ $=x^{3}+(-43t^{4}+86t^{3}+11t^{2}-38t+5)x^{2}$
$+(496t^{8}-1984t^{7}+1488t^{6}+1984t^{5}-1488t^{4}-992t^{3})x.$
There are four independent sections giving a discriminant of $4/3$ relative to the height pairing.
Proof.
The reader is referred to [21]. ∎
Since the torsion subgroup has order $2$, this implies a Picard discriminant of $6\cdot 4\cdot 4\cdot 2\cdot 4/3\cdot 1/4=64$ up to sign. The generators are all defined over ${\mathbb{Q}}(\sqrt{5})$. In a few hours we can enumerate the $191$ definite lattices in the genus that are the frame lattices for elliptic fibrations on the surface.
We are able to exhibit a model of this surface in $\mathbb{P}^{5}$ with $6$ ordinary double points and no more. There do exist fibrations defined over ${\mathbb{Q}}$ with no section, but for historical reasons we use one that merely has no ${\mathbb{Q}}$-rational section.
Proposition 4.45.
There is a fibration with no ${\mathbb{Q}}$-rational section with reducible fibre types $I_{1}^{*},I_{0}^{*},I_{2}^{3}$ and is such that in the $\mathbb{P}^{5}$ model the $I_{0}^{*}$ fibre contains a conic.
The equation is a bit too complicated to show here, but we easily find the general fibre, pass to the Jacobian, and write down some sections. With some effort we find a model of this new K3 surface as $S_{6}\subset\mathbb{P}^{6}$ with one $A_{2}$ and three $A_{1}$ singularities.
Proposition 4.46.
The dimension of the Galois-fixed subgroup of the Picard group of $S_{6}$ is $14$, and $S_{6}$ has a fibration defined over ${\mathbb{Q}}$ with multisection degree $2$ over ${\mathbb{Q}}$ and over $\bar{\mathbb{Q}}$. The reducible fibres of the Jacobian are of type $I_{4}^{*},I_{0}^{*},I_{4}$, and the torsion subgroup has order $2$. Its Weierstrass equation is
$y^{2}=x^{3}+(1/4t^{3}+65/4t^{2}+1724t)x^{2}+(31t^{4}-4960t^{3}+198400t^{2})x.$
There is a section of height $1$.
The fibres of this fibration have degree $8$; the corresponding map to $\mathbb{P}^{1}$ and the Jacobian of the general fibre can be found without too much difficulty.
Thus we have a surface of Picard discriminant $-16$; however, it is not the same lattice as that of the Kummer surface of a product. It is difficult to find a good model of this surface; we make do with $S_{8}\subset\mathbb{P}^{8}$ whose singularities are of type $A_{4},A_{3},A_{1},A_{1},A_{1}$. Although the full Picard group is still defined only over ${\mathbb{Q}}(\sqrt{5})$, the subgroup defined over ${\mathbb{Q}}$ now has rank $16$.
Proposition 4.47.
There is a fibration on $S_{8}$ defined over ${\mathbb{Q}}$ whose Jacobian has fibres of types $III^{*},III^{*},I_{2},I_{2}$ and that has $2$-torsion. The Weierstrass equation is
$y^{2}=x^{3}+2180x^{2}+\frac{-16/5t^{2}+167872t-19681280}{t}x.$
The $III^{*}$ fibres are rational, so the existence of a section forces all of their components to be rational. The $I_{2}$ fibres, however, are still conjugate over ${\mathbb{Q}}(\sqrt{5})$, so the rational Picard rank is $17$. (Note that the geometric Mordell-Weil rank is 00 for this fibration.)
Before comparing to the Kummer surface, we pass to yet another K3 surface, this time with Picard discriminant $1$.
Proposition 4.48.
On $S_{8}$ there is a fibration with two $II^{*}$ fibres and trivial Mordell-Weil group. The Jacobian is isomorphic to
$C_{E_{8}}:y^{2}=x^{3}+545x^{2}+74272x+(1/25t^{2}+69760t+98406400)/t.$
Proof.
Finding the Picard class $C$ of such a fibration and expressing it in terms of the sections and vertical curves of the previous fibration is not very difficult. However, using our methods as described above it would be painful to find equations for the fibration; the small Picard discriminant makes it difficult or impossible to find a model in a projective space of low dimension with singularities of small degree. Instead we proceed as follows. Spreading out the equation above to a surface in $\mathbb{P}^{2}\times\mathbb{P}^{1}$ and placing it into $\mathbb{P}^{5}$ by the Segre embedding, we find a model with two $A_{6}$ and two $A_{1}$ singularities. We write down the two reducible fibres as subschemes $F_{1},F_{2}$ of this surface. Since $2H-C$ is effective, the fibration can be defined by the equations $(p_{1}:p_{2})$, where $p_{i}$ is a polynomial of degree $2$ vanishing on $F_{i}$. This constrains each of the $p_{i}$ to lie in a $3$-dimensional projective space. We now reduce mod $7$ and test possibilities until we find one whose base scheme has the correct degree, then verify that it is in fact an elliptic fibration with two $II^{*}$ fibres. It turned out to be easy to recognize the $p_{i}$ as reductions of polynomials from characteristic 00 and thus to find the equations of the fibration. ∎
The following proposition facilitates the comparison with the Kummer surface.
Proposition 4.49.
The surface with generic fibre $C_{E_{8}}$ also admits a fibration with an $I_{12}^{*}$ fibre whose Weierstrass equation is
$y^{2}=x^{3}+(8t^{3}-271t^{2}-32t+1022)x^{2}+961x.$
Proof.
This can be proved by a trick similar to, but simpler than, that of Proposition 4.48. Letting $x_{0},x_{1},x_{2}$ and $y_{0},y_{1}$ be the coordinates on $\mathbb{P}^{2}\times\mathbb{P}^{1}$, we find that $(x_{0}:x_{2})$ is also a genus-$1$ fibration on the spreading-out of $C_{E_{8}}$ to a surface in $\mathbb{P}^{2}\times\mathbb{P}^{1}$. (This was found from the Segre embedding by considering the linear system $H-2E$, where $E$ is the exceptional divisor at one of the $A_{8}$ singularities of the embedding.) It is now routine to find the equation, and by translating the $2$-torsion point to the origin and scaling away some obvious powers of $2$ we reach the form indicated above. ∎
The fact that this fibration has an $I_{12}^{*}$ fibre means that over ${\mathbb{C}}$ it is connected with Kummer surfaces by means of a $2$-isogeny known as a Shioda-Inose structure [22, Section 6].
For the rest of this section let $\alpha=\frac{1+\sqrt{5}}{2}$. We will verify that the fibration is isomorphic to one on a K3 surface that is dominated by the Kummer surface $\mathcal{K}$ of $E\times E^{\sigma}$, where $E:y^{2}+xy+\alpha y=x^{3}+(\alpha+1)x^{2}+(\alpha-5)x+(3\alpha-5)$ is an elliptic curve of conductor ${\mathfrak{p}}_{31}$ (namely 2.2.5.1/31.1-a/5). Indeed, this curve is isomorphic to $y^{2}=x(x-1)(x-\frac{8\alpha+3}{16})$, so we may apply our general formulas to find a model for the Kummer surface over ${\mathbb{Q}}$ (taking care as usual to use the correct twist). There is a fibration with two $III^{*}$ fibres and no section, but since the $III^{*}$ fibres are not defined over ${\mathbb{Q}}$ this model cannot be directly compared with the one found as an image of the Hilbert modular surface.
Proposition 4.50.
There is a fibration on $\mathcal{K}$ with an $I_{6}^{*}$ fibre and six of type $I_{2}$. It has sections defined over ${\mathbb{Q}}(\sqrt{5})$ but none defined over ${\mathbb{Q}}$.
Using a model of the Kummer surface in $\mathbb{P}^{4}$ with an $A_{3}$ and nine $A_{1}$ singularities, we are able to write an equation for the general fibre, which has degree $12$ in $\mathbb{P}^{4}$. The only difficulty is that projection from the worst singularity is a map of degree $2$; however, it is manageable to find the divisor map given by the degree-$4$ place supported there. This gives a model in the standard form of an intersection of two quadrics in $\mathbb{P}^{3}$. From this we wish to construct an elliptic surface with an $I_{12}^{*}$ fibre; that is only a matter of taking an isogeny by the correct $2$-torsion point. We obtain the equation
$y^{2}=x^{3}+(8t^{3}-271t^{2}-32t+1022)x^{2}+961x,$
the same as in Proposition 4.49. Thus, in this case, we have proved the strongest form of the Oda-Hamahata conjecture.
5. Surfaces of general type
We now proceed to discuss the Hilbert modular surfaces of general type with $p_{g}=1$. For some of these, the Atkin-Lehner quotient is of general type and we can do nothing. However, there are other cases in which the quotient is of Kodaira dimension 00 or $1$. We prove a lemma to explain the difference.
Lemma 5.1.
Let $S$ be a smooth surface and $\phi:S\to S$ an involution whose fixed points are isolated and smooth, with $\phi$ acting on the tangent space at every fixed point by $-1$. Let $T=S/\phi$. Then $T$ has canonical singularities and $K_{S}=2\phi^{*}(K_{T})$.
Proof.
The singularities of $T$ occur only under the fixed points, and are locally isomorphic to $\mathbb{A}^{2}/\pm 1$, which is an $A_{1}$ singularity. For the statement on canonical divisors, let $S^{\prime}$ be the blowup of $S$ at the fixed points and $T^{\prime}$ the minimal resolution of $T$. Let $E_{1},\dots,E_{n}$ be the exceptional divisors of $S^{\prime}$. We then have $K_{S^{\prime}}=K_{S}+\sum_{i=1}^{n}E_{i}$ and $K_{T^{\prime}}=K_{T}$. In addition, the map $S^{\prime}\to T^{\prime}$ is ramified along the $E_{i}$, so the Riemann-Hurwitz formula gives $K_{S^{\prime}}=\phi^{*}(K_{T^{\prime}})+\sum_{i=1}^{n}E_{i}$. The result follows. ∎
Corollary 5.2.
With notation as above we have $K_{S}^{2}=2K_{T}^{2}$.
Proof.
This follows immediately from the push-pull formula: $K_{S}^{2}=2K_{S}\phi^{*}\cdot K_{T}=2\phi_{*}K_{S}\cdot K_{T}=2K_{T}^{2}$. ∎
Example 5.3.
One standard example of this situation arises when $S$ is an abelian surface and $\phi$ is the negation map.
A converse to Lemma 5.1 can hold only under additional assumptions: if the fixed points are singular, that does not imply that the quotient is not of general type. However, if they are not canonical, we expect that $2K_{T}^{2}<K_{S}^{2}$, which may imply that $K_{T}^{2}\leq 0$. If $K_{S}^{2}=2$, this implies that $K_{T}^{2}\leq 0$, and if $T$ is minimal that means that $\kappa(T)<2$. We will study such examples in Sections 5.1, 5.2, 5.3, proving the Conjecture in the last two cases.
Remark 5.4.
Some of the surfaces of general type have interesting special properties. As one example, we point out that the ring of modular forms for level $3{\mathfrak{p}}_{5}$ over ${\mathbb{Q}}(\sqrt{5})$ is generated by $5$ forms of weight $2$ and one of weight $4$, and that it is a module of rank $2$ over the subring $R$ generated by the forms of weight $2$. As a result the Hilbert modular variety in this case admits an additional automorphism (not one of the Atkin-Lehner involutions) arising from the double cover of the surface $\operatorname{Proj}(R)$. Though this is not the only example, we expect that such additional automorphisms are rare, just as in the situation of $X_{0}(N)$ where the only non-modular automorphism of a curve of general type occurs for $N=37$. Unfortunately the quotient by automorphisms acting as $+1$ on $\operatorname{H}^{2,0}$ still appears to be of general type.
The remaining cases occur at level ${\mathfrak{p}}_{17},{\mathfrak{p}}_{23}$ over ${\mathbb{Q}}(\sqrt{13})$, which both give surfaces of general type that are difficult to analyze, and level $(6)$ over ${\mathbb{Q}}(\sqrt{5})$, where I have not been able to obtain a usable form of the ring of modular forms from [2].
5.1. Level $2{\mathfrak{p}}_{7}$ over ${\mathbb{Q}}(\sqrt{2})$
The Kodaira dimension of this surface is left open by Hamahata; however, the software package determines that it has a model with $K^{2}=4$. Since $\operatorname{h}^{2,0}=1$ for this surface, it is of general type. Although we are not able to prove the Oda-Hamahata conjecture in this case, we can describe some interesting properties of the surface.
The surface has $6$ cusps corresponding to the divisors of $2{\mathfrak{p}}_{7}$. The Atkin-Lehner involution corresponding to ${\mathfrak{p}}_{7}$ acts on the set of cusps without fixed points, so the quotient is still of general type. On the other hand, the involution corresponding to $(2)$ fixes the cusps corresponding to $\sqrt{2}$-adic valuation $1$.
Since the Atkin-Lehner involutions act as $+1$ on the eigenform of weight $2$, the quotient still has $\operatorname{h}^{2,0}=1$. We will show that the quotient admits a fibration by curves of genus $1$.
Remark 5.5.
Computationally, this example is convenient in some ways, since the ring of modular forms is generated by the six Eisenstein series and one cusp form, all in weight $2$. In fact, the Hilbert modular surface is just a complete intersection of four quadrics in $\mathbb{P}^{6}$, although this misleadingly suggests that the canonical divisor would be ${\mathcal{O}}(1)$. The cusp singularities are not canonical, and in fact the canonical divisors are precisely the sections of ${\mathcal{O}}(1)$ that vanish on them, which constitute a $1$-dimensional space (as expected, since this is the dimension of the space of cusp forms). Taking the quotient by the Atkin-Lehner involutions, we obtain another intersection of four quadrics in $\mathbb{P}^{6}$, but this one is more singular and so its canonical divisor is smaller.
We will show that the Atkin-Lehner quotient of the Hilbert modular surface $H_{(2(3-\sqrt{2})),{\mathbb{Q}}(\sqrt{2})}$ admits a map to $\mathbb{P}^{1}$ whose generic fibre is of genus $1$ and that has double and triple fibres. Thus the minimum section degree is a multiple of $6$ (in fact exactly $6$), which makes the surface difficult to work with because Magma does not have a function to compute the Jacobian of a curve of genus $1$ unless there is a multisection of degree at most $5$. In order to analyze the surface properly, we therefore begin by studying the quotient only by the Atkin-Lehner operator at $2$. This will be a double cover of the full quotient unramified at the general fibre, and so it still has Kodaira dimension $1$. For this surface, which we call $\mathcal{E}_{3}$, the general fibre admits a $3$-section. It is possible (although not easy) to determine $\mathcal{E}_{3}$ as an elliptic surface with an explicit map to $\mathbb{P}^{1}$ and isomorphism of its generic fibre to a plane cubic. We state the result as a theorem.
Theorem 5.6.
$\mathcal{E}_{3}$ is an elliptic surface with two triple fibres and a $3$-section. Its Jacobian is isomorphic to the rational elliptic surface $y^{2}=x^{3}+(t^{2}+3)x^{2}+(2t^{2}+t+1/8)x$.
Proof.
The main steps involved in this calculation are as follows:
  1. (1)
    Having already written down the Hilbert modular surface in $\mathbb{P}^{6}$, we start by writing down the map to $\mathbb{P}^{12}$ given by the sections invariant under the Atkin-Lehner involution at $(2)$ and finding its image as a surface of degree $32$ in $\mathbb{P}^{12}$.
  2. (2)
    This surface has four obvious singularities, namely the images of the cusps. Projecting away from these produces a surface in $\mathbb{P}^{8}$ of degree $24$ singular along four lines, which are in bijection with the cusps of the Atkin-Lehner quotient.
  3. (3)
    We consider the linear system of quadrics vanishing twice on three of these lines and once on the other one (the choice does matter computationally; we single out the line corresponding to the cusp 00). This linear system defines a map to a surface of degree $14$ in $\mathbb{P}^{6}$, defined by $3,8,2$ equations of degree $2,3,4$ respectively.
  4. (4)
    Although this surface has bad singularities, a canonical sheaf computation in Magma yields a useful result: in particular, the divisor map associated to this sheaf has codomain $\mathbb{P}^{1}$ and generic fibre of degree $6$ and arithmetic genus $1$.
  5. (5)
    There is a curve on the surface that gives a $3$-section of the fibration (it happens to be isomorphic to an elliptic curve of conductor $14$ with $6$ rational torsion points). Projecting the generic fibre away from the intersection gives a model of degree $3$ in $\mathbb{P}^{3}$, and removing the linear equation gives a plane cubic.
  6. (6)
    Simplifying the model by moving some special fibres to lie above small points of $\mathbb{P}^{1}$ and then using general procedures for minimizing and reducing equations of elliptic surfaces gives the model above.
Having done this computation, we find that there are $3$ values of $t$ for which the plane cubic becomes a triple line; one is rational and the other two are conjugate over ${\mathbb{Q}}(\sqrt{-3})$. The rational one, however, is reduced in the model in $\mathbb{P}^{6}$ and is therefore not a multiple fibre of the surface. ∎
Remark 5.7.
The $j$-invariant of the given Weierstrass equation is invariant under the automorphism of ${\mathbb{Q}}(t)$ taking $t\to(12-t)/(1+4t)$, and in fact the surface is isomorphic to its base change by this map. This is as expected, since we know that there is an Atkin-Lehner involution acting on the surface. The fixed fibres are at $t=-2,3/2$.
We find that, in addition to the two triple fibres of type ${}_{3}I_{0}$, the elliptic surface has two fibres of type $I_{4}$ and four of type $I_{1}$. Letting the reduced subschemes of the triple fibres be $R_{1},R_{2}$, we use Kodaira’s canonical bundle formula [3, Theorem V.12.1] to compute that the canonical bundle is $2R_{1}+2R_{2}-F$, where $F$ is the class of a fibre. Thus the dimension of the linear system $|nK|$ is $\lfloor\frac{n}{3}\rfloor+a(n)$, where $a(n)=1,0,1$ for $n\equiv 0,1,2\bmod 3$ respectively.
We now return to the full Atkin-Lehner quotient. This allows us to determine $\mathcal{E}_{6}$ rigorously, even though we cannot do the computation directly in characteristic 00. Namely:
Theorem 5.8.
$\mathcal{E}_{6}$ is a surface with a genus-$1$ fibration and a double and a triple fibre. In suitable coordinates, its Jacobian has $j$-invariant $\frac{64t^{6}-1152t^{4}+6912t^{2}-13824}{t^{2}-8}$ and its singular fibres are $I_{4}$ at infinity, $I_{1}$ at ±$\pm\sqrt{8}$, and $I_{0}^{*}$ at $3$.
Proof.
The Atkin-Lehner operator at ${\mathfrak{p}}_{7}$ descends to an involution on $\mathcal{E}_{3}$, which exchanges the two triple fibres. It also exchanges the singular reduced fibres in pairs, so the quotient has one $I_{4}$ and two $I_{1}$ fibres. The two fixed points on the base may give a double fibre or an $I_{0}^{*}$ fibre, depending on whether the action there is a translation or a negation or a negation map. The Euler characteristic of the quotient must be a multiple of $12$, so exactly one of them must be an $I_{0}^{*}$ fibre, and the canonical bundle of $\mathcal{E}_{6}$ cannot be more positive than that of $\mathcal{E}_{3}$, so the other must be a double fibre.
One calculates that the Atkin-Lehner involution on the model above has $4$ fixed points; in the coordinates above, they lie on the fibre above $-2$, so that is the one that becomes an $I_{0}^{*}$. Since the fibre above $3/2$ is not fixed pointwise, the automorphism must be given by translation there. This specifies it uniquely, since there is only one rational $2$-torsion section. (In terms of the model above, the generic fibre is isomorphic to its base change by $t\to(12-t)/(4t+1)$ by an automorphism taking $(x:y:z)$ to $(x/((4t+1)/7)^{2}:y/((4t+1)/7)^{3}:z)$, which acts as the identity on the fibre at $3/2$ and as $-1$ on the fibre at $-2$. This automorphism must be composed with a translation, since the Atkin-Lehner involution does not fix a fibre pointwise, and that translation can only be by the $2$-torsion section.)
To find the $j$-invariant, we express the $j$-invariant of $\mathcal{E}_{3}$ in terms of the invariant function $(12-t)/(1+4t)+t$. The singular fibres are found by evaluating the coordinates of the singular fibres and the fibre on which the Atkin-Lehner automorphism acts by translation, as described above. ∎
Remark 5.9.
This is a member of the family with the given singular fibres that is the first entry of [15, Table 3].
Remark 5.10.
One could try to do the calculation directly by considering forms of degree $2$ vanishing on the worst singularities; this gives a map to a surface of degree $32$ in $\mathbb{P}^{12}$ that is singular along two lines. Projecting away from these we find a surface in $\mathbb{P}^{8}$ of degree $20$, and projecting away from two more obvious singularities gives yet another surface of degree $16$ in $\mathbb{P}^{6}$, which we will denote by $\mathcal{E}_{6}$. For the reduction of $\mathcal{E}_{6}$ mod small primes, Magma is able to compute, not merely the canonical divisor, but its tensor powers, and we find the dimension of the space of global sections of $nK$ to be $0,1,1,1,1,2$ for $1\leq n\leq 6$, and that the $6$-canonical map is a genus-$1$ fibration. Such a surface cannot have multisection degree less than $6$, and indeed we can find a $6$-section. However, it is difficult to do this calculation rigorously in characteristic 00.
The quotient by the product of the two Atkin-Lehner operators still has $\operatorname{h}^{2,0}=1$ and is of general type. It is a double cover of $\mathcal{E}_{6}$ branched along a $12$-section of self-intersection $-4$. Unfortunately, we are not able to exhibit a correspondence between this and a K3 surface. We leave this problem for future work.
5.2. Level $4{\mathfrak{p}}_{5}$ over ${\mathbb{Q}}(\sqrt{5})$
This surface is similar in some ways to the surface considered in Section 5.1, since as in that example the level is the product of the square of a prime and a prime. However, there is an important difference in that the eigenvalues for the weight-$2$ cusp form are $+1$ for all Atkin-Lehner operators, so the fact that the quotient is not of general type allows us to prove the Oda-Hamahata conjecture in this case. The group of Atkin-Lehner operators has order $4$ and there are $6$ cusps. The ring of Hilbert modular forms is generated by $7$ forms of weight $2$ and $3$ of weight $4$. The Atkin-Lehner at $4$ fixes two of the cusps, which allows the quotient not to be of general type, in light of Lemma 5.1.
In this case we will content ourselves with exhibiting a correspondence over ${\mathbb{Q}}(\zeta_{20})^{+}$. To be precise:
Theorem 5.11.
Let $S$ be the Hilbert modular surface $H_{4{\mathfrak{p}}_{5},{\mathbb{Q}}(\sqrt{5})}$ and let $K$ be the Kummer surface of $E\times E^{\prime}$, where $E,E^{\prime}$ are the elliptic curves 2.2.5.1/80.1-a/5 and 2.2.5.1/80.1-a/7. Then there is a dominant map $S\to K$ defined over ${\mathbb{Q}}(\zeta_{20})^{+}$, and in particular there is a correspondence embedding the transcendental lattice of $K$ into that of $S$.
Since these two curves are conjugate over ${\mathbb{Q}}$, we expect that such a correspondence exists even over ${\mathbb{Q}}$, but we have not shown this. We note that this argument is considerably easier than some of the previous ones even though we are working with a surface of general type.
Proof.
As usual, we start by mapping to $\mathbb{P}^{7}$ by forms of degree $2$ invariant under the involutions, obtaining a surface of degree $16$. Projecting away from the two cusps gives a surface of degree $10$, and although this is not actually a canonical model Magma’s CanonicalSheaf command produces a divisor class for which the associated map is a genus-$1$ fibration. The reducible fibre types are $I_{1}^{*},I_{0}^{*},I_{0}^{*},I_{4}$, and there is a $2$-torsion point; the Weierstrass equation is
$y^{2}$ $=x^{3}+(-2t^{4}+32t^{3}+172t^{2}-1168t+1254)x^{2}+$
$\qquad(t^{8}+32t^{7}+276t^{6}-368t^{5}-10050t^{4}+6848t^{3}+126916t^{2}-295152t+184041)x.$
Passing to the quotient gives us a fibration with fibres $I_{2}^{*},I_{0}^{*},I_{0}^{*},I_{2},I_{2}$ and full level-$2$ structure, and by a further $2$-isogeny we reach a surface whose reducible fibres are $I_{4}^{*},I_{0}^{*},I_{0}^{*}$ and that has a $2$-torsion section. This is already a type of fibration that occurs on the Kummer surface of a product of two elliptic curves, and so we can identify it using the formulas of [19, 4.2].
Over ${\mathbb{Q}}(\zeta_{20})^{+}$, the elliptic curves 2.2.5.1/80.1-a/5, 2.2.5.1/80.1-a/7 have full level-$2$ structure, so we see that the elliptic curves are quadratic twists of $y^{2}=x(x-1)(x-\lambda_{i})$ for
$\lambda_{1}=20\alpha^{3}+32\alpha^{2}-32\alpha-39,\quad 28\alpha^{3}-32\alpha^{2}-104\alpha+121$
by $-3\alpha^{2}+2\alpha+7$ and $82\alpha^{3}+3\alpha^{2}-6\alpha-$ 8 respectively, where $\alpha=\zeta_{20}+\zeta_{20}^{-1}$. Substituting into the formula for a fibration with $I_{4}^{*}$ and two $I_{0}^{*}$ fibres and changing coordinates on the base, we check that this one is isomorphic to the image of the Hilbert modular surface as described above. ∎
Remark 5.12.
The argument above cannot be trivially modified to produce a correspondence over ${\mathbb{Q}}$, since it seems not to be possible to define the appropriate fibration on the Kummer surface over ${\mathbb{Q}}$. It may be possible to do so over ${\mathbb{Q}}(\sqrt{5})$, however.
5.3. Level ${\mathfrak{p}}_{5}{\mathfrak{p}}_{11}$ over ${\mathbb{Q}}(\sqrt{5})$
We consider the Hilbert modular surface of level ${\mathfrak{p}}_{5}{\mathfrak{p}}_{11}$ over ${\mathbb{Q}}(\sqrt{5})$. Since $5$ is ramified, this is well-defined up to conjugacy. The level is not Galois stable, so the conjecture was not previously known. We will show that the Hilbert modular surface admits a dominant map to the Kummer surface of $E\times E^{\prime}$ over ${\mathbb{Q}}({\sqrt{5}})$, where $E,E^{\prime}$ are the elliptic curves 2.2.5.1/55.1-a/4, 2.2.5.1/55.2-a/6.
As usual, we start by using the software to compute the ring of modular forms and the Atkin-Lehner involutions. The standard model has $K^{2}=4$, as noted; since the fixed points of the Atkin-Lehner involutions are isolated and do not include cusps, by Corollary 5.2 one might expect the quotient to have $K^{2}=1$ and thus to be of general type. However, we will see that this is not the case.
The reason for this is that the elliptic points are among the fixed points of $w_{{\mathfrak{p}}_{5}}$. The software computes that there are $4$ elliptic points, two each of types $(5;1,2)$ and $(5;1,3)$. These are not canonical singularities (a cyclic quotient singularity is canonical if and only if its type is of the form $(n;d,n-d)$ for some $d<n$). They are in fact the points where all modular forms of weight $2$ and $4$ vanish, and one computes that they are fixed by $w_{{\mathfrak{p}}_{5}}$. Thus Lemma 5.1 does not apply.
Proposition 5.13.
The full Atkin-Lehner quotient is birational to a K3 surface $S_{5}$ in $\mathbb{P}^{5}$ that has $12$ ordinary double points and no other singularities. These points comprise one orbit defined over ${\mathbb{Q}}(\sqrt{5},\sqrt{11})$, one over ${\mathbb{Q}}(\sqrt{5},i)$, and two over ${\mathbb{Q}}(\sqrt{5})$.
Proof.
Embedding the quotient by ${\mathcal{O}}(2)$, we find a surface of degree $16$ in $\mathbb{P}^{7}$; projecting away from the cusp, we have a surface singular along a line. The projection has degree $7$ in $\mathbb{P}^{4}$ and is a nonminimal surface which is the blowup of a K3 surface at a smooth point. Magma readily finds the minimal model and verifies that the singularities are as claimed. ∎
Proposition 5.14.
There is a genus-$1$ fibration on $S_{5}$ whose reducible fibres are of type $I_{1}^{*},I_{0}^{*},I_{3},I_{3}$ and whose Mordell-Weil group is of rank $3$ and defined over ${\mathbb{Q}}(\sqrt{5})$.
Since the expected Picard rank of the K3 surface in this case is $18$, these are all of the generators (and this can be checked by verifying that the zeta function of the reduction mod $19$ has an irreducible factor of degree $4$). The Mordell-Weil lattice has discriminant $1$, indicating that the Picard lattice discriminant is $144=16\cdot 3^{2}$, and one checks that it is contained in the Picard lattice of the Kummer surface of a product with index $3$. The entire Picard lattice is defined over ${\mathbb{Q}}(\sqrt{5})$, and the rational sublattice has rank $13$. Unfortunately there does not seem to be a fibration defined over ${\mathbb{Q}}$ with multisection degree $3$.
Proposition 5.15.
There is a genus-$1$ fibration defined over ${\mathbb{Q}}$ on $S_{5}$ with a section over ${\mathbb{Q}}(\sqrt{5})$, with multisection degree $2$, and with reducible fibres $I_{3}^{*},I_{0}$, and three $I_{2}$. The rational sublattice of the Picard lattice of the Jacobian has rank $16$.
This is done by computation. As usual, the reader is referred to [21] for the details; the equation of the general fibre is too large to appear here.
Definition 5.16.
Let $S^{\prime}$ be the Jacobian of the fibration of Proposition 5.15.
We again attempt to find a fibration on $S^{\prime}$ defined over ${\mathbb{Q}}$ with multisection degree $3$, but without success. There are rational fibrations with multisection degree $2$. However, computing equations for such a fibration, we found that the Jacobian did not appear to have any fibrations defined over ${\mathbb{Q}}$ with no rational section, or any that admit an isogeny that would reduce the discriminant, and so we were unable to proceed. Instead we will work over ${\mathbb{Q}}(\sqrt{5})$.
Proposition 5.17.
There is a genus-$1$ fibration on $S^{\prime}$ defined over ${\mathbb{Q}}({\sqrt{5}})$ with multisection degree $3$ whose Jacobian is isomorphic to the Kummer surface of $E\times E^{\prime}$ over ${\mathbb{Q}}({\sqrt{5}})$, where $E,E^{\prime}$ are the elliptic curves 2.2.5.1/55.1-a/4, 2.2.5.1/55.2-a/6.
Proof.
Again, this is proved by computation. We find a good model of $S^{\prime}$ in $\mathbb{P}^{7}$ and study the curves on it and the ADE configurations that they form, eventually discovering a fibration with multisection degree $3$ and two $I_{2}^{*}$ and four $I_{2}$ fibres. This is the same as one of the configurations on the Kummer surface, so we attempt to match them using the formulas for fibrations from [20]. This is successful with one of the $9$ types of ${I_{2}^{*}}^{2}I_{2}^{4}$ fibration for the indicated elliptic curves. ∎
The following summarizes the contents of this section.
Theorem 5.18.
The Oda-Hamahata conjecture holds for elliptic curves over ${\mathbb{Q}}(\sqrt{5})$ with discriminant $\sqrt{5}(4-\sqrt{5})$, in the sense that the desired correspondence exists over ${\mathbb{Q}}(\sqrt{5})$.
5.4. Level ${\mathfrak{p}}_{13}$ over ${\mathbb{Q}}(\sqrt{3})$
For our final example, we briefly discuss a case in which the field has narrow class number $2$ rather than $1$ and the geometric genus is $2$: namely, we choose the level to be a prime ideal of norm $13$ in ${\mathbb{Z}}[\sqrt{3}]$. (I thank Damián Gvirtz-Chen for the suggestion to study such examples.) The Hilbert modular surface has two components and an Atkin-Lehner operator of level ${\mathfrak{p}}_{13}$ acts on their union. Since ${\mathfrak{p}}_{13}$ has a totally positive generator $4+\sqrt{3}$, the Atkin-Lehner operator preserves the components rather than exchanging them, and there is no reason for these to be isomorphic.
Proposition 5.19.
Both components admit dominant rational maps to K3 surfaces of Picard rank $18$, the positive component of discriminant $-196$ and the negative component of discriminant 784=4196-784=4$\cdot$-196- 784 = 4 ⋅ - 196. These components have the same number of ${\mathbb{F}}_{p}$-points for $p<100$ of good reduction.
Proof.
This is a straightforward application of techniques used many times already. After several steps we find a map from the positive component to a K3 surface of Picard rank $18$ and discriminant $-196$ (it has a fibration with reducible fibres of type $I_{2},I_{4},I_{4},I_{6}$ and Mordell-Weil lattice of rank $4$ and discriminant $49/3$). On the other hand, the negative component requires more work, as our initial constructions of elliptic fibrations have very large rank, which would make it difficult to find a Mordell-Weil basis. By finding additional fibrations not defined over ${\mathbb{Q}}$ we are, however, able to find a basis for the Picard lattice. ∎
Remark 5.20.
The Picard group of the negative component is defined over a large number field, which would make it difficult to compute directly with the curves on the surface. However, by reducing modulo a suitable prime we do not change the Picard lattice (since the rank is even) and this calculation is much easier. Since the Picard discriminant is a multiple of $7^{2}$ we do not expect to be able to complete the computation and do not attempt to determine the Galois action on the Picard lattice.
The Hilbert modular form of level ${\mathfrak{p}}_{13}$ appears to be connected with an elliptic curve $E/{\mathbb{Q}}(\zeta_{12})$ which is a “${\mathbb{Q}}(\sqrt{3})$-curve”, i.e., it is $2$-isogenous to its Galois conjugate over ${\mathbb{Q}}(\sqrt{3})$. The curve $E$ is defined by the equation
$y^{2}+(-\zeta_{12}^{2}-\zeta_{12}+1)xy+(z_{12}^{3}+z_{12}^{2})y=x^{3}+(z_{12}^{3}+z_{12}^{2})x^{2}$
and has conductor ${\mathfrak{p}}_{13}{\mathbb{Z}}[\zeta_{12}]$. To be precise about the connection, fix a prime ${\mathfrak{p}}$ of ${\mathbb{Z}}[\sqrt{3}]$, let $a_{\mathfrak{p}}$ be the trace of Frobenius for $E\bmod{\mathfrak{p}}$, and let $e_{\mathfrak{p}}$ be the Hecke eigenvalue for the Hilbert modular form. If ${\mathfrak{p}}$ is inert in ${\mathbb{Q}}(\zeta_{12})/{\mathbb{Q}}(\sqrt{3})$, then $a_{\mathfrak{p}}=e_{\mathfrak{p}}^{2}-N({\mathfrak{p}})$; otherwise $a_{\mathfrak{p}}=e_{\mathfrak{p}}$.
The torsion subgroup of $E$ has order $14$, which is presumably related to the factor of $7^{2}$ in the discriminant. This factor makes it difficult to relate the two K3 surfaces to each other or to the Kummer surface of $E\times E^{\sigma}$, but we certainly expect such relations to exist. In particular, we might ask whether there is a correspondence between the K3 surfaces arising from the positive and negative quotients; as above, there is no Atkin-Lehner operator relating them, so there is no obvious way to construct one.
Just as ${\mathbb{Q}}$-curves over totally real fields correspond to base change Hilbert modular forms, these calculations suggest a generalization of Hamahata’s conjecture to certain ${\mathbb{Q}}(\alpha)$-curves:
Conjecture 5.21.
Let $f$ be a Hilbert cuspidal newform of level $I$ over the field $K={\mathbb{Q}}(\sqrt{D})$ such that there exist a quadratic field ${\mathbb{Q}}(\sqrt{n})$ and a quadratic extension $L/K$ such that $e_{\mathfrak{p}}(f)$ is rational for ${\mathfrak{p}}$ split in $L/K$ and $\sqrt{n}$ times a rational number for ${\mathfrak{p}}$ inert in $L/K$. Then there exist an elliptic curve $E/L$ which is isogenous to its Galois conjugate over $K$, a K3 surface $S$ defined over ${\mathbb{Q}}$ and $K$-isogenous to the Kummer surface of $E\times E^{\sigma}$ and admitting a correspondence to the Hilbert modular surface $H_{I,K}$, where $I$ is the conductor of $E$ viewed as an ${\mathcal{O}}_{K}$-ideal.

### v2

@@
 Let $E$ be an elliptic curve over a real quadratic field belonging to an isogeny class with one of the following LMFDB labels or their Galois conjugates: 2.2.5.1/31.1-a, 2.2.5.1/55.1-a, 2.2.5.1/80.1-a, 2.2.8.1/9.1-a, 2.2.8.1/17.1-a, 2.2.8.1/32.1-a, 2.2.13.1/9.1-a, 2.2.17.1/4.1-a, 2.2.17.1/8.3-a.
-Then there is a correspondence (not necessarily defined over ${\mathbb{Q}}$) between the Hilbert modular variety of level equal to the conductor of $E$ and the Kummer surface of $E\times E^{\sigma}$.
+Then there is a correspondence (not necessarily defined over ${\mathbb{Q}}$) between the Hilbert modular variety of level equal to the conductor of $E$ and the Kummer surface of $E\times E^{\sigma}$, inducing a map on $\operatorname{H}^{2}$ that is injective on the transcendental lattice of the Kummer surface.
 To give some context for this result, let us more generally consider an elliptic curve $E$ over a totally real number field $K$.
@@
 Let $E$ be such a curve and let its conductor be $I$.
-To generalize the algebraic statement of modularity is fairly straightforward: one requires that the coefficients of the $L$-function of $E$ match the Hecke eigenvalues of a Hilbert modular form $f$.
+To generalize the algebraic statement of modularity is fairly straightforward.
+One requires that the coefficients of the $L$-function of $E$ match the Hecke eigenvalues of a Hilbert modular form $f$.
 This is known for all elliptic curves over real quadratic fields and many over fields of higher degree: see [10, Theorems 1-2].
@@
 194] Let $E$ be an elliptic curve over a real quadratic field $K$ with conductor $I$.
-Then there is a correspondence between $H_{I,K}$ and $K_{E}=(E\times E^{\sigma})/\pm 1$ that induces an injection from the primitive part of $\mathcal{H}^{2}(K_{E})$ to that of $\mathcal{H}^{2}(H_{I,K})$.
+Then there is a correspondence between $H_{I,K}$ and $K_{E}=(E\times E^{\sigma})/\pm 1$ that induces an injection from the primitive part of $\operatorname{H}^{2}(K_{E})$ to that of $\operatorname{H}^{2}(H_{I,K})$.
 Remark 1.10.
@@
 An especially strong form of this conjecture would state that this correspondence can be defined over ${\mathbb{Q}}$, not just over $\bar{\mathbb{Q}}$.
+Remark 1.11.
+Hamahata does not define the term “primitive”, but we understand the phrase “primitive part of $\operatorname{H}^{2}$” to refer to the transcendental lattice, namely the quotient of $\operatorname{H}^{2}$ by the subspace spanned by fundamental classes of curves.
+We also remind the reader that a correspondence $X\subset V\times W$ induces maps on cohomology [37, p.
+3].
+We have $\operatorname{H}^{*}(V\times W)\cong\operatorname{H}^{*}(V)\otimes\operatorname{H}^{*}(W)$ by the Künneth formula, and $\operatorname{H}^{*}(V)\cong\operatorname{H}^{*}(V)^{\vee}$ by Poincaré duality.
+Thus $\operatorname{H}^{*}(V\times W)\cong\operatorname{Hom}(\operatorname{H}^{*}(V),\operatorname{H}^{*}(W))$, and one sees easily that if $X\subset V\times W$ with $\dim X=\dim V=\dim W$ (the only case used here) the map preserves degrees.
+Our correspondences will not be explicit subvarieties of $V\times W$; rather, they are induced from pairs of rational maps $\phi:V\dashrightarrow T,\psi:W\dashrightarrow T$ as the closure of the locus of $(v,w)$ where $\phi(v),\psi(w)$ are both defined and are equal.
 Remark 1.12.
@@
 This is not clearly stated or proved in the literature, but is known to experts.
-It is a consequence of the modularity of $E$, which is proved in [27, Corollary 6.2], subject to Serre’s conjecture on modular $2\times 2$ Galois representations, which is now a theorem of Khare and Wintenberger [17] More precisely, Ribet’s result is that $E$ is a quotient of $J_{1}(N)$; it follows that there is a dominant map $J_{1}(N)^{d}\to\prod_{i=1}^{d}E^{\sigma_{i}}$, where $d=[K:{\mathbb{Q}}]$.
+It is a consequence of the modularity of $E$, which is proved in [27, Corollary 6.2], subject to Serre’s conjecture on modular $2\times 2$ Galois representations, which is now a theorem of Khare and Wintenberger [17].
+More precisely, Ribet’s result is that $E$ is a quotient of $J_{1}(N)$; it follows that there is a dominant map $J_{1}(N)^{d}\to\prod_{i=1}^{d}E^{\sigma_{i}}$, where $d=[K:{\mathbb{Q}}]$.
 The domain of this map is related to $\operatorname{H}^{d}(X_{0}(N))$ by Hecke correspondences, and the equality of the levels follows from Serre’s formula for the level of the modular form, also proved by Khare and Wintenberger.
@@
 In order to multiply two such expansions, one needs an algorithm for writing an element of ${\mathcal{O}}^{+}$ as a sum of two such elements in all possible ways.
-An efficient algorithm for doing so will be described in [1].
+An efficient method for doing so will be described in [1].
 We state our main result here:
@@
 In this case we do find a K3 surface that is in correspondence with the Hilbert modular variety (and in fact this has been known for some years, being [35, Theorem 9]) but it is quite difficult to establish an explicit correspondence with the desired Kummer surface.
-In two of the cases where $I=(1)$ the birational equivalence between the Hilbert modular surface and a K3 surface with canonical singularities is made explicit in [37, Appendix].
+In two of the cases where $I=(1)$ the birational equivalence between the Hilbert modular surface and a K3 surface with canonical singularities is made explicit in [38, Appendix].
 We have also considered a few cases in which the field $K$ has narrow class number $2$. One of these will be discussed in Section 5.4.
@@
 In some cases the Kummer surface of $E\times E^{\sigma}$ is itself a quotient of $H_{I,K}$; in other cases this does not appear to be true and instead we find an auxiliary K3 surface $S$ that admits maps of finite degree from both $H_{I,K}$ and the Kummer surface of $E\times E^{\sigma}$.
-The methods are similar to those of [37, Appendix], except that we do not necessarily start with a variety birational to a K3 surface.
+The methods are similar to those of [38, Appendix], except that we do not necessarily start with a variety birational to a K3 surface.
 In addition, the setup of [38] requires information about the ring of Hilbert modular forms that is not available here: in particular, that it can be generated by Borcherds products.
-Against this, we do not find equations for the Hirzebruch-Zagier cycles [36, Def.
+Against this, we do not find equations for Hirzebruch-Zagier cycles [36, Def.
 VI.1.4] on the Hilbert modular surfaces.
@@
 It is known [23, Theorem 1] that if a surface $S$ is of geometric genus $1$ then there is a K3 surface $K_{S}$ and a correspondence between $S$ and $K_{S}$ identifying the transcendental part of $H^{2}_{{\mathrm{\acute{e}t}}}$ for the two surfaces.
-However, finding such a K3 surface explicitly is a difficult and interesting problem in general.
+However, finding the K3 surface and correspondence explicitly is a difficult and interesting problem in general.
 In some such cases the quotient by some or all of the Atkin-Lehner involutions is a surface with $p_{g}=1$ and $\kappa<2$ and we can find the desired correspondence, but otherwise there is no obvious method to use.
@@
 The image $S_{L}$ of $S$ under projection away from $L$ is in a space of dimension $2$ less and satisfies $\deg S-\deg S_{L}\geq 5$, so such an operation always brings us closer to the degree for a standardly embedded K3 surface.
-It can happen that we run out of room, in the sense that projection away from a singular point or line gives a map of degree greater than $1$.
+The projection away from a singular point or line may be a map of degree greater than $1$.
 This is undesirable, because the image may no longer have the desired cohomology.
@@
 For almost all primes $p$, the reduction of $V$ mod $p$ will have the same property, and we can find the point mod $p$ by enumerating the ${\mathbb{F}}_{p}$-points of the reduction and checking them for singularity (deciding whether a point is singular is a simple computation if we know the dimension of $V$).
-By looking for rational numbers that reduce mod the various $p$ to the coordinates of these points, we guess the coordinates of the point over ${\mathbb{Q}}$.
+By looking for rational numbers that reduce modulo the various $p$ to the coordinates of these points, we guess the coordinates of the point over ${\mathbb{Q}}$.
 Again, this is not a proof, but we may easily check whether our candidate is actually on $V$ and, if so, whether it is singular.
@@
 Otherwise we may have to consider the reduced 00-dimensional schemes supported at the singular points mod $p$ and interpolate the coefficients in the Gröbner basis of the ideals defining them.
-If the points are not rational, then of course for some primes we will only be able to find their reductions by passing to an extension of ${\mathbb{F}}_{p}$, which is undesirable.
+If the points are not rational, then for some primes we will only be able to find their reductions by passing to an extension of ${\mathbb{F}}_{p}$, which is undesirable.
 Fortunately in this work it was rarely necessary to find singular points not defined over a quadratic field.
@@
 It may be defined by polynomials with unreasonably large coefficients, causing future computations to be slow.
-This situation can often be remedied as follows: find a minimal integral model and let $D(t)$ be the discriminant, ignoring any contribution at infinity so that this is a polynomial.
+This situation can often be remedied as follows: find a minimal integral model and let $D(t)=\prod_{i}t-\alpha_{i}$, where the $\alpha_{i}$ are the coordinates of singular fibres excluding $\infty$.
 Find an automorphism of $\mathbb{P}^{1}$ under which the numerator of the pullback of $D$ has smaller coefficients.
@@
 Up to Galois conjugacy, Hamahata [14] lists $14,6,4,2$ such surfaces over the real quadratic fields of discriminant $5,8,13,17$ respectively.
-(Such surfaces also exist over fields whose strict class number is not $1$, and over the fields of discriminant $29,37,41$ at level $1$, but these are not his concern.) We summarize his results on the Kodaira dimensions in the following table:
+(Such surfaces also exist over fields whose strict class number is not $1$, and over the fields of discriminant $29,37,41$ at level $1$, but these are not his concern.) We summarize his results on the Kodaira dimensions.
+Table 1. Kodaira dimensions of Hilbert modular surfaces with geometric genus $1$ over real quadratic fields with narrow class number $1$.
 $\kappa=0$ $\kappa=1$ $\kappa=2$ $\kappa\in\{1,2\}$
@@
 We intend to study these two situations in more detail in future work.
-Of the $7$ cases with $p_{g}=1$ where the surface was known not to be of general type, there are only $2$ where the level is not Galois-stable and thus the conjecture is not already known.
+Of the $7$ cases with $p_{g}=1$ where the surface was known not to be of general type, there are only $2$ that do not correspond to ${\mathbb{Q}}$-curves, so that the conjecture is not already known.
 As mentioned above, the surface $H_{{\mathfrak{p}}_{31},{\mathbb{Q}}(\sqrt{5})}$ has Kodaira dimension $1$, and we will show that $H_{2{\mathfrak{p}}_{7},{\mathbb{Q}}(\sqrt{2})}$, another surface left undecided by Hamahata, is of general type.
@@
 Remark 4.9.
+The generic fibre has no ${\mathbb{Q}}(t)$-rational divisors of degree between 00 and $4$, because there are fibres with no points over any quadratic extension of ${\mathbb{Q}}_{3}$.
 We do not know whether the generic fibre admits any ¯$\bar{\mathbb{Q}}(t)$-rational divisors of degree between 00 and $4$, but we suspect not.
@@
 Namely, we consider the isogeny class 2.2.13.1/9.1-a of elliptic curves of conductor $(3)$ over ${\mathbb{Q}}(\sqrt{13})$.
-Again, in this case the conductor is Galois-stable and so the Oda-Hamahata conjecture is already known.
+Again, we have a ${\mathbb{Q}}$-curve and so the Oda-Hamahata conjecture is already known.
 Nevertheless, the explicit calculation is of some interest.
@@
 4.5. Level $(2)$ over ${\mathbb{Q}}(\sqrt{17})$
-In this section we consider a case where the level is Galois stable and thus the conjecture is known. Nevertheless there are some interesting questions to study.
+In this section we consider a ${\mathbb{Q}}$-curve, for which the conjecture is already known. Nevertheless there are some interesting questions to study.
 Since $2$ splits in ${\mathbb{Q}}(\sqrt{17})$, there are $4$ cusps and hence $4$ Eisenstein series in weight $2$.
@@
 In fact the Picard lattice is isomorphic to that of a K3 surface with $4$ fibres of type $I_{5}$, a section of height $8/5=4-2(6/5)$, and a $5$-torsion section.
-The $5$-isogeny goes to a surface of Picard discriminant $200$; however, as before it seems impossible to find a genus-$1$ fibration on this surface defined over ${\mathbb{Q}}$ with multisection degree $5$, so we cannot find the correspondence with a Kummer surface over ${\mathbb{Q}}$.
+The $5$-isogeny goes to a surface of Picard discriminant $200$; however, as in Remark 4.30 it seems impossible to find a genus-$1$ fibration on this surface defined over ${\mathbb{Q}}$ with multisection degree $5$, so we cannot find the correspondence with a Kummer surface over ${\mathbb{Q}}$.
 We proceed to find a correspondence with a Kummer surface over ${\mathbb{Q}}(\sqrt{2},\sqrt{3})$, the $2$-division field of two of the curves in the isogeny class.
@@
 The minimal desingularization of $S_{5}$ has the same Hodge diamond as a K3 surface, but it is not itself a K3 surface.
-This follows, for example, from Kodaira’s formula for the canonical bundle of a surface with a genus-$1$ fibration, combined with the existence of a point of $\mathbb{P}^{1}$ where the fibre is nonreduced on $S_{5}$ but is a smooth rational curve on the Jacobian.
+This follows from Kodaira’s formula for the canonical bundle of a surface with a genus-$1$ fibration, combined with the existence of a point of $\mathbb{P}^{1}$ where the fibre is nonreduced on $S_{5}$ but is a smooth rational curve on the Jacobian.
 The situation is analogous to that of an Enriques surface, for which the Hodge diamond is the same as that of a rational elliptic surface.
@@
 We will verify that the fibration is isomorphic to one on a K3 surface that is dominated by the Kummer surface $\mathcal{K}$ of $E\times E^{\sigma}$, where $E:y^{2}+xy+\alpha y=x^{3}+(\alpha+1)x^{2}+(\alpha-5)x+(3\alpha-5)$ is an elliptic curve of conductor ${\mathfrak{p}}_{31}$ (namely 2.2.5.1/31.1-a/5).
-Indeed, this curve is isomorphic to $y^{2}=x(x-1)(x-\frac{8\alpha+3}{16})$, so we may apply our general formulas to find a model for the Kummer surface over ${\mathbb{Q}}$ (taking care as usual to use the correct twist).
+Indeed, this curve is isomorphic to $y^{2}=x(x-1)(x-\frac{8\alpha+3}{16})$, so we may apply our general formulas to find a model for the Kummer surface over ${\mathbb{Q}}$, taking care as usual to use the correct twist.
 There is a fibration with two $III^{*}$ fibres and no section, but since the $III^{*}$ fibres are not defined over ${\mathbb{Q}}$ this model cannot be directly compared with the one found as an image of the Hilbert modular surface.
@@
 The curve $E$ is defined by the equation
-$y^{2}+(-\zeta_{12}^{2}-\zeta_{12}+1)xy+(z_{12}^{3}+z_{12}^{2})y=x^{3}+(z_{12}^{3}+z_{12}^{2})x^{2}$
+$y^{2}+(-\zeta_{12}^{2}-\zeta_{12}+1)xy+(\zeta_{12}^{3}+\zeta_{12}^{2})y=x^{3}+(\zeta_{12}^{3}+\zeta_{12}^{2})x^{2}$
 and has conductor ${\mathfrak{p}}_{13}{\mathbb{Z}}[\zeta_{12}]$.